### v1

# Physics-Driven Construction of Compact Primitive Gaussian Density Fitting Basis Sets

Abstract
We present model-assisted density fitting (MADF) basis set generator, an algorithm for generating primitive atomic Gaussian density fitting (DF) basis sets (DFBSs) from a contracted Gaussian orbital basis set (OBS). The MADF algorithm produces DFBSs suitable for accurate robust DF approximation of 2-particle interactions in mean-field and correlated electronic structure. The algorithm is designed to (a) saturate the OBS product space by a large regularized set of primitive solid-harmonic Gaussian shells with nonuniform distribution of exponents followed by (b) pruning of the shells according to their contributions to the 2-body energy of a correlated atomic ensemble. Building the DFBS generator model almost exclusively on mathematical and physical principles allows one to limit the number of parameters that control the density fitting error to four, with a single set of parameters sufficient for computations with all basis cardinal numbers, with and without correlation of core electrons, with and without scalar and spin-dependent relativistic effects, spanning almost all of the Periodic Table. Performance assessment included basis sets up to quadruple-zeta quality from several major basis set families, using molecules composed of main-group, d-block, and f-block elements. The resulting DF errors in Hartree-Fock and second-order MP2 energies (with relativistic all-electron treatments, when appropriate) were on the order of $20$ and $10$ $\mu E_{h}$ per electron, respectively.
1 Introduction
Density fitting (DF)1, 2, 3, 4, 5, 6, also known as resolution of identity (RI),7, 8 is a widely used cost and complexity reduction technique for approximate evaluation of various operators in the electronic structure theory, most commonly the electron repulsion integrals (ERIs). The cost of ERI evaluation and their storage are often crucial limitations in practical simulations of electronic structure. The factorization of ERIs using the DF approximation can be exploited in various ways to gain significant speedups and reduce storage requirements9, 10, 11, 12, 13, 14, 15, 16, such as to reduce the cost of electrostatic potential evaluation10, 17 and to facilitate correlated and relativistic computations17, 18, 19, 20, 21. DF has also recently emerged as the route to deeper factorizations of the ERI tensor such as real-space tensor hypercontraction,22 algebraic pseudospectral,23 canonical polyadic,24 and interpolative separable density fitting (ISDF)25. Although our sole focus in this work is on the Gaussian AOs, note that density fitting is a key enabling technology for other choices of AOs such as Slater-type26 and numerical AOs.27
In the DF approach, the AO products (or “densities”) in the ERIs are expanded in a special-purpose density-fitting (or auxiliary) basis of AOs. Traditionally, density-fitting basis sets (DFBSs) are manually-optimized for each orbital basis set (OBS) in ad hoc manner28, 29, 17, 30, 31, 32, 33, 34, 35, 36, 37. For efficiency reasons not only are DFBS matched to the OBS but they are also matched to specific use cases; e.g., DFBS for Coulomb fitting (RI-J)29, 31, Coulomb and exchange fitting (RI-JK)17, 32 or for the second-order Møller-Plesset (MP2) correlation energy (RI-C)28, 17, 30, 33, 34, 36, 37. DFBS optimization is usually done with cumbersome iterative HF and MP2 calculations on a training set of atoms and molecules.
Due to the significant effort involved in the DFBS development, there are only a few OBS families for which there are matching DFBS, and those often only focus on the top half of the Periodic Table. Among the commonly used basis sets, only the correlation consistent basis sets and the def2- (TURBOMOLE) basis sets have matching DFBS. Even for these families there are signficant coverage gaps; e.g., correlation-consistent OBSs such as cc-pVXZ, cc-pCVXZ, and cc-pwCVXZ38, 39, 40 have DFBS coverage gaps in the first to third rows of the Periodic Table.41 Only the def2- family has a thorough DFBS coverage for atoms up to Rn (with atoms heavier than Kr using effective core potential).17, 31, 41, 42, 43, 44 Unfortunately, for heavier elements, there are no useful DFBS coverage. For example, OBS families such as ANO-RCC basis sets45, 46, 47, cc-pwCVXZ-DK / -DK3 / -X2C48, 49, 50, 51 basis sets, the x2c-XZVP52, 53 family of basis sets, and Dyall basis sets54, 55, 56 do not have the corresponding DFBS designed for correlation calculations. The only two relativistic DFBSs available in Basis set exchange (BSE)41 database are x2c-JFIT and x2c-universal-JFIT53, but they are designed only for Coulomb fitting and are not suitable for many-body correlated simulations or even for hybrid Kohn-Sham DFT computations.
To address the challenges of ad hoc optimization of DFBS many black-box methods have been proposed for generating (rather than optimizing) DFBS for a given OBS have been proposed57, 58, 59, 60, 61, 62, 63 . A DFBS generator can be viewed as an OBS $\to$ DFBS map controlled by zero or more model parameters. The key difference between DFBS generators and manual optimization is that no nonlinear optimization is involved in the definition of the map. A map can be defined straightforwardly to produce exact atomic DFBS (= DFBS that is exact for any computation on a single atom), but such DFBS is too large to be practical. Nevertheless, such an exact map is typically the starting point for the design of approximate DFBS generators. Nearly all DFBS generators in practice are approximate, i.e. they produce DFBS that are not exact even for a single atom.
Most DBFS generators are designed only for DF in the context of evaluating the Coulomb potential of a charge density. These include the AutoABS method by Yang et al.57, long-range corrected auxiliary basis sets by Hellmann et al.58, even tempered auxiliary basis sets with shared exponents by Daz-Tinoco et al.59 and a similar approach has also recently been reported for relativistic OBSs60.
DFBS generators designed for mean-field and correlated electronic structure simulation include the AutoAux procedure by Stoychev et al.,61 methods proposed by Lehtola62, 63, and PySCF built-in generator of even-tempered DFBSs.
The AutoAux method, available in the ORCA software package64 and in the command-line interface of Basis Set Exchange (BSE)41, produces a primitive even-tempered DFBS that can be applied universally (in mean-field and correlated simulations). The AutoAux generator uses even-tempered sequences of Gaussian primitive AOs to span the OBS-deduced range of exponents for each angular momentum block. The exponent ranges and even-tempered ratios vary with the angular momenta ($L$) and the atomic number $Z$ (these even-tempering parameters are prescribed for up to $L=7$; the BSE implementation of AutoAux uses the same exponent ratio for $L\geq 7$). The generator also limits the highest angular momentum of DFBS AOs below what is needed for exact atomic DF, in a manner that depends on the highest angular momentum of the occupied orbitals. AutoAux as a result is quite complex, defined with more than a dozen model parameters in total.
PySCF’s simple built-in generator65 produces even-tempered primitive DFBS and is very similar in spirit to AutoAux. It has fewer model parameters (e.g., even-tempering ratio is not graded with angular momenta) and claims to offer similar accuracy to AutoAux, although no extensive comparison exists.
Lehtola presented a DFBS generator that produces nearly exact atomic DFBS composed of primitive solid-harmonic Gaussian (SHG) AOs.62 First, the OBS AOs on a given atom are used to generate a base pool of primitive SHGs by approximating each angular momentum channel of solid-harmonic OBS AO products by a single solid-harmonic AO. The resulting pool is then regularized by pivoted Cholesky decomposition (pCD) of the two-center two-electron Coulomb integral matrix, with the accuracy of the resulting primitive DFBS controlled robustly by a single threshold. Although the resulting DFBSs are nearly exact, they are extremely large, lead to high condition numbers in molecules, and contain primitive Gaussians with twice the angular momentum of the highest angular momentum ($L_{\text{OBS}}$) of functions in the parent OBS. Lehtola addressed the latter issue in his recent work66 by extending the algorithm to prune high-angular momentum functions out of the DFBS with a tunable parameter. The algorithm then uses Kállay’s67 contraction method with singular value decomposition (SVD) to contract the DFBS, further reducing its size.
An attractive alternative to density fitting that avoids these issues and provides some of its advantages is the (pivoted) Cholesky decomposition (CD).68 Any positive-definite 2-particle interaction represented in an arbitrary OBS can be decomposed with accuracy controlled by a single threshold. CD has been used for electronic structure simulations in both relativistic69, 70, 71 and nonrelativistic72, 73, 74 contexts. Although the accuracy of CD can be robustly pushed beyond the reach of conventional density fitting, the cost of CD is substantial. Computing analytic derivatives of the Cholesky vectors is also fundamentally challenging. The 1-center CD (1C-CD),75 later renamed atomic CD (aCD),76 works around the problem of analytic derivatives of CD by marrying CD with density fitting. Namely, global CD is avoided and CD is only used to construct a contracted DFBS for each atom suitable for accurate representation of a given operator in a given OBS. Thus the aCD approach can be viewed as a black-box 1-parameter DFBS generator. The aCD approach also produces contracted DFBSs which have mixed angular momentum functions and mixed representations (cartesian or spherical). A solution to this issue was proposed by Hellmann et al.77, combining aCD and Lehtola’s method62, however, it suffers from the same issues as Lehtola’s original method.62 Several challenges in designing efficient CD algorithms in the context of quantum chemistry have also been mentioned.78
To navigate the severe gaps in the coverage by manually-constructed DFBS of existing work and the significant shortcomings of the existing DFBS generators we attempted to design a new DFBS generator that produces (a) primitive Gaussian DFBSs, (b) of comparable size and accuracy as the manually-optimized DFBS, (c) usable in the context of mean-field and correlated simulation, and (d) has as few adjustable parameters as possible to ensure maximum universality. By combining the existing ideas for spanning the product space of OBS AOs with new ideas for regularizing and pruning the exact DFBS using atomic correlated ensemble RDMs we arrived at a DFBS generator dubbed MADF. The purpose of this manuscript is to describe it and assess its performance. Following a quick recap of density fitting in Section 2.1 we describe the 2 key ingredients of the MADF generator: the regularization of the complete set of candidate primitives (described in Section 2.2) and subsequent pruning by a 2-body energy estimator model (described in Section 2.3). Section 3 describes the technical and details of the parameter training regiment. In Section 4 we document the performance of the DFBSs generated by our method against the manually-optimized DFBSs and the DFBSs generated by the AutoAux algorithm.
2 Formalism
2.1 Density Fitting
DF approximates product of AOs (“density”) $\phi_{\mu}({\bf r})\phi_{\nu}({\bf r})\equiv({\bf r}|\mu\nu)$ as a linear combination of DFBS AOs $\phi_{X}({\bf r})$
$|\mu\nu)\approx\sum_{X}C_{\mu\nu}^{X}|X)\equiv|\widetilde{\mu\nu}).$ (1)
The fitting coefficients $C_{\mu\nu}^{X}$ are obtained by minimizing a norm of the density error $|\delta_{\mu\nu})\equiv|\mu\nu)-|\widetilde{\mu\nu})$. To minimize the error in the diagonal elements of the matrix representation of a positive operator $\hat{O}$, namely in $(\mu\nu|\hat{O}|\mu\nu)$, the optimal choice of the error norm is the expectation value of operator $\hat{O}$ itself:
$||f||_{\hat{O}}\equiv(f|\hat{O}|f),$ (2)
This is known as robust density fitting.79, 80 Least-squares minimization of $||\delta_{\mu\nu}||_{\hat{O}}$ is equivalent to solving the following equation
$\hat{O}|\mu\nu)=\sum_{Y}\hat{O}|Y)C^{Y}_{\mu\nu}.$ (3)
In global DF Eq. 3 is projected on every DF AO in the system to produce the usual system defining the fitting coefficients:
$(X|\hat{O}|\mu\nu)=\sum_{Y}(X|\hat{O}|Y)C^{Y}_{\mu\nu}.$ (4)
Clearly, DF is a misnomer (and so is RI), as in the most common scenario, what is being “fitted” (minimized) is the $L^{2}$ norm of $\hat{O}^{1/2}|\delta_{\mu\nu})$, not of $|\delta_{\mu\nu})$ itself. For the most important case where $\hat{O}$ is the Coulomb potential operator $\hat{J}$,
$({\bf r}|\hat{J}|f)\equiv\int\mathrm{d}\mathbf{r}^{\prime}\,|{\bf r}-{\bf r}^{\prime}|^{-1}f({\bf r^{\prime}}),$ (5)
the robust fitting minimizes the $L^{2}$ norm of the fitting error in the electric field generated by the AO product.
DFBS generators usually57, 58, 59, 60, 61, 62, 63, 64 e.g., AutoAux, aCD work in a single-atom context, i.e., DFBS is constructed using information about products on OBS AOs on a single atom. It is reasonable to question whether this is sufficient. Assuming a system composed of a single atom type, global density fitting in a system of $n$ atoms will be able to fit every such 1-center product of OBS AOs using $n$ times as many DFBS AOs as in a single atom. Indeed, for long-range $\hat{O}$ the fitting of even strongly-localized AO products generally involves DFBS AOs from far away.81 Thus, 1-center products in a molecule can borrow DFBS AOs from other centers to make the fitting more accurate than in a single atom. This suggests that a single atom poses a more challenging DF setting than a molecule, which supports the use of a single-atom setting for the generation of DFBS.
2.2 Products of primitive Gaussians AOs
The density fitting becomes exact if the span of DFBS matches the span of the OBS AO products. Because DFBS AOs are atom-centered but the products of AOs in a molecule are generally centered between atoms, exact DF is only possible for a single atom. We then confine ourselves to a single atom. Furthermore, we will limit ourselves to the case of OBS and DFBS composed of solid-harmonic Gaussian (SHG) AOs. Unnormalized primitive SHG with exponent $\alpha$ is defined as:82, 83
$\chi_{\alpha,l,m}({\bf r})=$ $r^{l}\exp(-\alpha r)y_{lm}(\theta,\phi)$ (6)
$y_{lm}(\theta,\phi)=$ $(-1)^{m}P_{l}^{|m|}(\theta)\begin{cases}\cos(m\phi),\quad m\geq 0\\ \sin(|m|\phi),\quad m<0\end{cases}.$ (7)
Whereas product of any number of concentric primitive Cartesian Gaussians is a single Cartesian Gaussian (modulo normalization), product of concentric SHGs $\chi_{\alpha_{\mu},l_{\mu},m_{\mu}}$ and $\chi_{\alpha_{\nu},l_{\nu},m_{\nu}}$ is a linear combination of $r^{l_{\mu}+l_{\nu}-L}\chi_{\alpha_{\mu}+\alpha_{\nu},L,m_{\mu}+m_{\nu}}$ for $\max(|l_{\mu}-l_{\nu}|,m_{\mu}+m_{\nu})\leq L\leq l_{\mu}+l_{\nu}$. Clearly, the contribution to the product from the Clebsch-Gordan channel $L<l_{\mu}+l_{\nu}$ is an SHG multiplied by $r^{l_{\mu}+l_{\nu}-L}$. Lehtola proposed to account for this effect by approximating each such contribution by a single SHG of angular momentum $L$ with the following effective exponent (see Appendix II of Ref. 62):
$\alpha_{\text{eff}}^{L}=\left[\frac{\Gamma(l_{\mu}+l_{\nu}+\frac{3}{2})\Gamma(L+2)}{\Gamma(L+\frac{3}{2})\Gamma(l_{\mu}+l_{\nu}+2)}\right]^{2}(\alpha_{\mu}+\alpha_{\nu})\begin{cases}=\alpha_{\mu}+\alpha_{\nu},\quad L=l_{\mu}+l_{\nu}\\ >\alpha_{\mu}+\alpha_{\nu},\quad L<l_{\mu}+l_{\nu}\end{cases}.$ (8)
We follow this prescription here to define the “complete” set of candidate primitive SHGs for a given product of primitive SHG OBS AOs.
Let $\mathcal{C}$ denote the “complete” set of all unique solid harmonic primitive AOs generated thereby from all products of primitive OBS AOs on a given atom. The set $\mathcal{C}$ can be used as DFBS for nearly exact density fitting in an atom (a) for any contraction of the generating set of primitive OBS AOs and (b) for any operator. However, even in a single atom such DFBS is highly linearly dependent and overcomplete, 84, 62 therefore, it is unusable with finite precision arithmetic. Moreover, such a basis would be extremely large, defeating the purpose of the DF approximation. Hence, a compact DFBS is needed to represent the product space of AO functions. Since the set $\mathcal{C}$ is overcomplete for use as a DFBS for an atom, the unimportant and linearly dependent functions must be trimmed out of this set. Lehtola proposed to do such pruning by pCD.62 However, even after pCD the resulting DFBSs are still large, as illustrated in Figure 1. While pCD reduces the density of the complete set, it largely leaves its range of exponents intact (at least for lower $L$), and it still keeps many pairs of adjacent exponents with small ratios ($\ll 2$). The problem is especially pronounced for the high angular momentum ($L=6$).
(a) OBS = def2-QZVPP
(b) OBS = x2c-QZVPPall-2c
Figure 1: Exponents of DFBS generated (or manually-optimized) for Kr atom with representative nonrelativistic and relativistic OBS using various methods scattered on a logarithmic number line for different angular momenta i.e., $L=\{0,1,2,6\}$. The number of exponents is shown on the right of each scatter plot. Figure 1(a) shows distribution of DFBS exponents for nonrelativistic OBS def2-QZVPP and Figure 1(b) for relativistic OBS x2c-QZVPPall-2c. The plot depict Lehtola’s complete set (shown in purple), Lehtola’s pCD-regularized set (shown in red) with pCD threshold $10^{-7}$ as recommended in Ref.62, even-tempered exponents produced by AutoAux (shown in green) and set of exponents produced by the MADF algorithm (shown in orange). Distribution of exponents of manually-optimized DFBS def2-QZVPP-RIFIT is also shown (in blue) in Figure 1(a).
The AutoAux exponent ranges are significantly reduced relative to that of Lehtola’s pCD-derived DFBS, and the exponents are well separated owing to even-tempering. However, it is clear that the optimal DFBS exponents are not necessarily even-tempered, as the exponent density of the complete and pCD-processed DFBS of Lehtola and that of the manually-optimized DFBS has significant nonuniformity. Ideally, one would also be able to avoid varying the exponent ratios with $L$ for the even-tempering recipe.
In this work, we regularize the complete candidate exponent pool using a simple adaptive algorithm designed to produce non-even-tempered sets that deviate from the input set as little as possible and with exponent ratios as least as large as the target ratio threshold. Given a primitive SHG set $\mathcal{C}$ and a target exponent ratio threshold $\zeta$, each subset $\mathcal{C}_{L}$ containing primitives of angular momenta $L$ is regularized by repeatedly replacing the pair of primitive SHG shells with the smallest exponent ratio $\alpha_{1}/\alpha_{2}$ by a single primitive SHG shell with the geometric mean exponent $\alpha=\sqrt{\alpha_{1}\alpha_{2}}$. Regularization stops if no pairs of primitives have an exponent ratio less than $\zeta$. Clearly, regularization is most easily implemented if $\mathcal{C}_{L}$ is first sorted by the exponent in descending order. The algorithm is described in detail in Algorithm 1. The recommended value of $\zeta$ will be determined heuristically in Section 4.1.
Algorithm 1 Generation of regularized complete set of DF SHG shells.
1:function RegularizedCompleteDFBS($\mathcal{P}$, $\zeta$) $\triangleright$ $\mathcal{P}$: primitive OBS shells; $\zeta\geq 1$
2:     $\mathcal{C}\leftarrow\textsc{CompleteDFBS}(\mathcal{P})$triangleright Complete set of DF SHG shells with exponents generated by Eq. 8
3:     $L_{\text{max}}\leftarrow\textsc{MaxAM}(\mathcal{c})$triangleright Maximum angular momentum of shells in $\mathcal{c}$
4:     ${}\mathcal{R}\leftarrow\{\}$triangleright Initialize the regularized complete DFBS
5:     for $L\in[0,L_{\text{max}}]$ do
6:         $\mathcal{c}\leftarrow\textsc{SelectByAM}(\mathcal{C},L)$triangleright Shells in $\mathcal{C}$ with angular momentum $L$
7:         if Size($\mathcal{c}$)$\geq 2$ then
8:              $\zeta_{\text{min}}\leftarrow\infty$triangleright smallest exponent ratio
9:              $I_{\text{min}}\leftarrow 0$
10:              while $\zeta_{\text{min}}<\zeta$ do
11:                  $\mathcal{c}\leftarrow\textsc{SortByExponent}(\mathcal{c})$triangleright Sort in order of decreasing exponents
12:                  for $i$ in $20\dots\textsc{Size}(\mathcal{c})-$ 2 do
13:                       $\zeta_{i}\leftarrow\textsc{Exponent}(\mathcal{c}\left[i\right])/\textsc{Exponent}(\mathcal{c}\left[i+1\right])$
14:                       if $\zeta_{i}<\zeta_{\text{min}}$ then
15:                           $\zeta_{\text{min}}\leftarrow\zeta_{i}$
16:                           $I_{\text{min}}\leftarrow i$triangleright Tracks shell pair with smallest exponent ratio
17:                       else
18:                           continue
19:                       end if
20:                  end for
21:                  $\alpha_{I}\leftarrow\textsc{Exponent}(\mathcal{c}\left[I_{\text{min}}\right])$
22:                  $\alpha_{I+1}\leftarrow\textsc{Exponent}(\mathcal{c}\left[I_{\text{min}}+1\right])$
23:                  $c\leftarrow\textsc{SHGShell}(\sqrt{\alpha_{I}\alpha_{I+1}},L)$triangleright Pair of shells with closest exponents fused to single shell
[…]
25:              end while
26:         end if
27:         $\mathcal{R}\leftarrow\textsc{Concat}(\mathcal{R},\mathcal{c})$
28:     end for
29:     return $\mathcal{R}$
30:end function
2.3 Estimating energy contribution of DFBS primitives
In a single atom, DFBS must contain solid harmonics with $L$ up to $2L_{\text{OBS}}$ to obtain exact values for all possible matrix elements of the Hamiltonian in OBS, and the above regularization methods do not change this fact. However, it is well-known that the importance of DF AOs with the highest angular momenta is relatively low, at least for energies. Manually-optimized DFBSs typically restrict the $L_{\text{DFBS}}$ to significantly below $2L_{\text{OBS}}$, since only DFBS with $L$ up to $L_{\text{OBS}}+L_{\text{occ}}$ are needed for exact representation of $(ov|ov)$ integrals needed for the MP2 energy. In practice MP2-targeted DFBS are used successfully in the context of higher-order methods, like coupled-cluster. In ad hoc DFBS generators like AutoAux, the angular momenta of DFBS are restricted heuristically, i.e. $L_{\text{max,DF}}=\text{max}(2L_{\text{occ}},L_{\text{OBS}}+L_{\text{inc}})$, where $L_{\text{inc}}$ is fixed to 1 or 2 depending on $Z$61 To avoid the heuristics as much as possible, we attempted to design a first-principles energy-based model for further pruning of DFBS.
Consider the expectation value of a 2-body Fock-space Hamiltonian:
$E=\sum_{pq}h_{p}^{q}\gamma_{q}^{p}+\frac{1}{2}\sum_{pqrs}h_{pq}^{rs}\gamma_{rs}^{pq},$ (9)
where $h_{p}^{q}$ and $h_{pq}^{rs}$ are the Hamiltonian matrix elements and $\gamma^{p}_{q}$ and $\gamma_{rs}^{pq}$ are the elements of the 1- and 2-body reduced density matrices, respectively. Since DF is used to approximate the Coulomb integrals $h_{pq}^{rs}\equiv(pr|qs)\equiv(pr|\hat{J}|qs)$, we will focus on the two-body part of the energy:
$^{2}E=\frac{1}{2}\sum_{pqrs}(pr|qs)\gamma_{rs}^{pq}.$ (10)
Our objective is to quantitatively estimate the energetic contribution of a given AO in a given DFBS to ${}^{2}E$, using a first-principles model for the 2-RDM. The most straightforward approach would be to evaluate ${}^{2}E$ with the Coulomb integral approximated via robust DF,
$\left(pr|qs\right)=\sum_{XY}(pr|X)({\bf J}^{-1})_{XY}(Y|qs),$ (11)
where $(pq|X)\equiv(pq|\hat{J}|X)$ and $({\bf J})_{XY}\equiv(X|\hat{J}|Y)$. The magnitude of the change in DF-approximated ${}^{2}E$ due to the inclusion of the candidate DF AO in DFBS could be used to gauge its importance for accurate computation of ${}^{2}E$. There are two problems with such an approach. positive (electrostatic) and negative (exchange, correlation) contributions. Thus, contributions from a given DF AO to the positive and negative parts of ${}^{2}E$ could cancel each other spuriously, causing an artificial omission of such AO from DFBS. Replacing the summands in Eq. 10 by their magnitudes would resolve the sign issue but would, in effect, change the weights of different components of the energy. Second, it is known that the charge density fitting used to compute the electrostatic energy only requires fitting functions with lower angular momenta; thus, it makes sense to tune the fitting basis to the more challenging exchange and correlation contributions to the energy. Therefore, we model the energetic importance of DF AOs by the diagonal exchange-like contribution to ${}^{2}E$ obtained from Eq. 10 by substitution $rs\to qp$:
$^{2}E^{\text{dx}}\equiv$ $\frac{1}{2}\sum_{pq}(pq|qp)\gamma_{qp}^{pq}=-\frac{1}{2}\sum_{pq}(pq|qp)\gamma_{pq}^{pq}.$ (12)
Note that $\gamma_{pq}^{pq}\geq 0$, hence each summand in Eq. 12 is positive (real orbitals are assumed for brevity).
Another question is how to obtain the 2-RDM. Our original efforts to model the importance of DFBS focused on heuristic models of RDMs (such as the Fermi-Dirac distribution). Thus in practice we constructed an estimator that requires a model for orbital occupancies $n_{p}\equiv\gamma_{p}^{p}$ only, rather than the full 2-RDM. To model the 2-RDM in terms of orbital occupancies, we use the Cauchy-Schwarz inequalities for the 2-RDM,82 which for positive $\gamma^{pq}_{pq}$ becomes simply:
$\gamma^{pq}_{pq}\leq\sqrt{n_{p}n_{q}}.$ (13)
This leads to the following bound for ${}^{2}E^{\text{dx}}$ in terms of orbital occupancies:
$|^{2}E^{\text{dx}}|\equiv-^{2}E^{\text{dx}}\leq$ $\frac{1}{2}\sum_{pq}(pq|qp)\sqrt{n_{p}n_{q}}\equiv~{}^{2}\bar{E}^{\text{dx}}.$ (14)
This expression is independent of whether the single-particle states are spin-free or spin-orbital. Since robust DF guarantees that the error in $(pq|qp)=(pq|pq)$ is positive:
[…]
then the robust DF approximation to Eq. 14,
$^{2}\tilde{\bar{E}}^{\text{dx}}\equiv$ $\frac{1}{2}\sum_{XY}\sum_{pq}(pq|X)({\bf J}^{-1})_{XY}(Y|pq)\sqrt{n_{p}n_{q}},$ (16)
is guaranteed to be smaller than the exact value. This gives us the ability to estimate the importance of a given DF AO for modeling ${}^{2}\bar{E}^{\text{dx}}$ by monitoring its effect on ${}^{2}\tilde{\bar{E}}^{\text{dx}}$ as well as the ability to control the overall error in ${}^{2}\bar{E}^{\text{dx}}$.
The angular momentum conservation ensures that $X$ and $Y$ in Eq. 16 must have the same angular momentum. This makes it convenient to split ${}^{2}\tilde{\bar{E}}^{\text{dx}}$ into its angular momentum components:
$^{2}\tilde{\bar{E}}^{\text{dx}}_{L}\equiv$ $\frac{1}{2}\sum_{\begin{subarray}{c}XY\\ l_{X}=l_{Y}=L\end{subarray}}\sum_{pq}(pq|X)({\bf J}^{-1})_{XY}(Y|pq)\sqrt{n_{p}n_{q}},$ (17)
where $l_{X}$ is the angular momentum of DF AO $X$. Practical tests showed that it is important to represent ${}^{2}\tilde{\bar{E}}^{\text{dx}}_{L}$ for $L\leq 2L_{\text{occ}}$ more accurately than the higher $L$ channels. Practical tests also showed that ${}^{2}\tilde{\bar{E}}^{\text{dx}}_{L}$ grows with $Z$, hence it makes sense to prune DFBS according to the value of ${}^{2}\tilde{\bar{E}}^{\text{dx}}_{L}$ per unit nuclear charge.
The next question is in which order to consider the candidate DF AOs. To decide the order we rewrite the ${}^{2}\tilde{\bar{E}}^{\text{dx}}_{L}$ in terms of contributions from individual DF AOs:
$^{2}\tilde{\bar{\epsilon}}^{\text{dx}}_{X}\equiv$ $\frac{1}{2}\sum_{\begin{subarray}{c}Y\\ l_{Y}=l_{X}\end{subarray}}\sum_{pq}(pq|X)({\bf J}^{-1})_{XY}(Y|pq)\sqrt{n_{p}n_{q}}$ (18)
${}^{2}\tilde{\bar{E}}^{\text{dx}}_{L}=$ $\sum_{\begin{subarray}{c}X\\ l_{X}=L\end{subarray}}{}^{2}\bar{\epsilon}^{\text{dx}}_{X}$ (19)
Although ${}^{2}\bar{\epsilon}^{\text{dx}}_{X}$ are not strictly positive, they are nearly so; all negative values are very small (on the order of tens of microhartrees). For each angular momentum channel $L$ vector $|^{2}\bar{\epsilon}^{\text{dx}}_{X}|$ is computed once and sorted in descending order; this defines the order in which DF AOs are considered.
2-body energy pruning of the candidate DF AO pool $\mathcal{R}^{L}$ produced by Algorithm 1 proceeds as follows. If ${}^{2}\tilde{\bar{E}}^{\text{dx}}_{L}$ from (Eq. 17) is less than $Z\tau$, where $\tau$ is the target threshold, then all DF AOs of this angular momentum are discarded, else the DF AOs are sorted according to the order of importance provided by $|^{2}\bar{\epsilon}^{\text{dx}}_{X}|$. The candidate DF AOs are added to the $L$-channel of pruned DFBS, $\mathcal{D}^{L}$, until the difference between ${}^{2}\tilde{\bar{E}}^{\text{dx}}_{L}$ evaluated with $\mathcal{R}^{L}$ and with $\mathcal{D}^{L}$ is smaller than $Z\tau$. For an atom with ground-state configuration including occupied subshells of angular momenta $\leq L_{\text{occ}}$ DF channels $L\leq 2L_{\text{occ}}$ are pruned with $\tau=\tau_{1}$, the rest are pruned with $\tau=\tau_{2}$. Using the same values of $\tau$ for $Z=1$ (hydrogen) as for other $Z$ produced larger errors than desired due to the overpruning of the higher $L$ DF channels; thus, separate threshold $\tau_{\text{H}}$ is used for pruning DF channels of all angular momenta for H. The values of $\tau_{1}$, $\tau_{2}$, and $\tau_{\text{H}}$ will be determined heuristically in Section 4.1.
2.4 Correlated model of orbital occupancies
Our initial instinct was to model orbital occupancies in Eq. 14 by Fermi-Dirac distribution (with heuristics for the temperature). Unfortunately such an approach did not prove fruitful. Using atomic mean-field ensemble orbitals often used as initial guess for atomic density matrix would not suffice, since this will fail to generate nonzero populations for orbitals important in dynamical electron correlation. Therefore we developed a simple model for correlated 1-RDM of a neutral atom in an ensemble state using spin-opposite scaled (SOS) first-order Møller-Plesset perturbation of a model mean-field ensemble density. The mean-field density in OBS is obtained by one-shot diagonalization of OBS Fock matrix $\mathbf{F}$ computed from the atomic ensemble density generated by populating subshells of a minimal basis using the atom’s reference ground-state electron configuration. Such definition for the OBS Fock matrix is utilized in the Libint library and the MPQC program under the name “superposition of atomic densities” (SOAD) for generating guess Fock matrices for SCF. SOAD can be viewed as a simplified version of the popular SAD method85, with differences described for completeness in appendix A.
Diagonalization of SOAD ensemble Fock matrix expressed in OBS AO (see appendix A),
$\mathbf{F}_{\mathrm{SOAD}}\mathbf{C}=\mathbf{SC}\bm{\varepsilon},$ (20)
produces energies and OBS AO coefficients of non-self-consistent ensemble orbitals. The orbitals are assigned spin-free occupancies $0\leq{\bf n}^{(0)}\leq 2$ using the same spherically-averaged ensemble definition used to construct the SOAD ensemble density and $\mathbf{F}_{\mathrm{SOAD}}$. Columns in ${\bf C}$ corresponding to zero and nonzero ensemble occupancies are deemed unoccupied ($a,b$) and occupied ($i,j$), respectively. These orbitals and the corresponding orbital energies are used to compute opposite-spin ensemble Møller-Plesset first-order amplitudes:
$t_{ab}^{ij}=\frac{\sqrt{n_{i}^{(0)}n_{j}^{(0)}}}{2}\frac{(ia|jb)}{\varepsilon_{i}+\varepsilon_{j}-\varepsilon_{a}-\varepsilon_{b}}.$ (21)
The corresponding second-order correlation contributions ${\bf n}^{(2)}$ to orbital occupancies are obtained straghtforwardly:
$n_{i}^{(2)}=$ $\sum_{jab}-2t^{ij}_{ab}t_{ij}^{ab},$ (22)
$n_{a}^{(2)}=$ $\sum_{ijb}2t^{ij}_{ab}t_{ij}^{ab}.$ (23)
The total correlated orbital occupancy vector ${\bf n}$ is a sum of the mean-field and correlated contributions:
${\bf n}=$ ${\bf n}^{(0)}+{\bf n}^{(2)}.$ (24)
2.5 The MADF Algorithm
For a given (input) OBS consisting of contracted SHG AOs on an atom with atomic number $Z$ MADF (1) generates a “complete” set of primitive DF SHG AOs necessary to represent all products of uncontracted OBS AOs using the method of Lehtola62, (2) regularizes it by fusing pairs of shells with overly close exponents as described in Section 2.2, and (3) prunes out DF AOs not important for description of correlated 2-body energy using the method described in Section 2.3. Complete description of the algorithm is given in Algorithm 2. Its 4 model parameters (with 5 unique values) summarized in Table 1.
Parameter Description Values
$\zeta$ Exponent ratio threshold for Algorithm 1 1.4
$\tau_{\text{H}}$ 2-body energy threshold for the hydrogen atom $10^{-6}$
$\tau_{1}$ 2-body energy threshold for $Z>1$ and $L\leq 2l^{occ}$ $10^{-6}$ or $10^{-8}$ [^1] Optimal values of $\tau_{1}$ are different for nonrelativistic and relativistic OBSs, see Sections 4.1.2 and 4.1.3
$\tau_{2}$ 2-body energy screening threshold for $Z>1$ and $L>2l^{occ}$ $10^{-5}$
Table 1: Model parameters of the MADF DFBS generator.
Algorithm 2 The MADF algorithm
1:function MADF($Z,\mathcal{O},\mathcal{M},\tau_{1},\tau_{2},\tau_{\text{H}},\zeta$) $\triangleright$ $Z$: atomic number; $\mathcal{O}$: (contracted) OBS; $\mathcal{M}$: MBS; $\tau_{\{1,2,\text{H}\}}\geq 0$; $\zeta\geq 1$
2:     $\mathcal{P}\leftarrow\textsc{Uncontract}(\mathcal{O})$
3:     $\mathcal{C}\leftarrow\textsc{RegularizedCompleteDFBS}(\mathcal{P},\zeta)$triangleright Algorithm 1
4:     $\mathbf{F}_{\mathrm{SOAD}}\leftarrow\textsc{SoadFock}(Z,\mathcal{O},\mathcal{M})$triangleright appendix A
5:     $\mathbf{F}_{\mathrm{SOAD}}\mathbf{C}=\mathbf{SC\epsilon}$triangleright Eq. 20
6:     ${\bf n}\leftarrow\textsc{CorrelatedOccupancies}(\mathbf{C},\mathcal{M})$triangleright Eqs. 21, 22 and 24
7:     $L_{\text{occ}}\leftarrow\textsc{MaxAM}(\mathcal{M})$triangleright maximum angular momentum of occupied subshells.
8:     $L_{\text{max}}\leftarrow\textsc{MaxAM}(\mathcal{R})$triangleright maximum angular momentum of candidate pool of shells
9:     $\mathcal{D}\leftarrow\{\}$triangleright Initialize the target list of DF shells
10:     for $L\in[0,L_{\text{max}}]$ do
11:         $\mathcal{d}\leftarrow\{\}$
12:         if $Z\neq 1$ then
13:              if $L\leq 2L_{\text{occ}}$ then
14:                  $\tau\leftarrow\tau_{1}$
15:              else
16:                  $\tau\leftarrow\tau_{2}$
17:              end if
18:         else
19:              $\tau\leftarrow\tau_{\text{H}}$
20:         end if
21:         $\mathcal{r}\leftarrow\textsc{SelectByAM}(\mathcal{R},L)$triangleright Shells in $\mathcal{R}$ with angular momentum $L$
22:         $\{E_{\mathcal{r}},\{\epsilon_{\mathcal{r}}\}\}\leftarrow\textsc{Eval2EdxL}(\mathcal{r})$triangleright Evaluate Eqs. 17 and 18 using $\mathcal{r}$ as DFBS
23:         if $E_{\mathcal{r}}\geq Z\tau$ then
24:              $\mathcal{r}\leftarrow\textsc{SortBy}(\mathcal{r},\{|\epsilon_{\mathcal{r}}|\})$triangleright Sort $\mathcal{r}$ in order of decreasing $|{}^{2}\tilde{\bar{\epsilon}}^{\text{dx}}_{X}|$
25:              $E_{\mathcal{d}}\leftarrow 0,i\leftarrow 0$
26:              while $E_{\mathcal{r}}-E_{\mathcal{d}}\geq Z\tau$ and $i<\textsc{Size}(r)$ do
27:                  $\mathcal{d}\leftarrow\textsc{Concat}(\mathcal{d},\mathcal{r}[i])$
28:                  $\{E_{\mathcal{d}},\_\}\leftarrow\textsc{Eval2EdxL}(\mathcal{d})$triangleright Evaluate Eq. 17 using $\mathcal{d}$ as DFBS
29:                  $i\leftarrow i+1$
30:              end while
31:              $\mathcal{D}\leftarrow\textsc{Concat}(\mathcal{D},\mathcal{d})$
32:         end if
33:     end for
34:     return $\mathcal{D}$
35:end function
3 Technical details
The MADF algorithm is implemented in a developmental version of MPQC86; generation of base DFBS by Lehtola’s algorithm62 was implemented in Libint.87, 88 For optimization of MADF parameters and its benchmark testing, we employed the nonrelativistic and relativistic Hartree-Fock (HF) and second-order Møller-Plesset (MP2) methods. DF errors were computed relative to the HF and MP2 energies obtained with exact (four-center) two-electron matrix elements throughout. The exact two-component decoupled one-body Hamiltonian89, 90, 91 (1eX2C, or simply X2C) with spin-orbit coupling was used for all relativistic computations without empirical scaling corrections. The core orbitals were frozen in nonrelativistic MP2 computations with cc-pVXZ and def2- OBSs.
The MADF model parameters for use with nonrelativistic OBS were determined on training set TS1 composed of 38 closed-shell systems with light elements: AlF3, Ar, BF3, BH3, C2H2, C2H4, C2H6, CH3OH, CH4, Cl2, CO2, COH2, CS2, F2, H2O, HCl, HF, HNO2, N2, cis-N2H2, trans-N2H2, Ne, NH3, NOCl, O2, O3, OF2, P2, PF3, PH3, S2, SF2, SiH4, SiO2, and SO2. The performance of the recommended model parameters was assessed on the standard G2 set of molecules92, containing 118 closed- and 30 open-shell systems. For training and assessment computations cc-pVXZ38, cc-pwCVXZ40(X = D, T, Q) and def2-SVP, def2-XZVP (X = T, Q)44 were used. Since there are no cc-pwCVXZ basis sets for Li, Be, and Na atoms, systems containing these atoms were skipped from the assessment of these OBSs.
The MADF model parameters for use with relativistic OBS were determined on training set TS2 composed of 24 closed-shell systems with heavy elements: AgH, AsH3, AuH, BiH3, CuH, GaH3, GeH4, HAt, HBr, Hi, InH3, Kr, PbH4, PoH2, PtC, RuC, SbH3, ScH, SeH2, SnH4, TeH2, TlH3, Xe and Rn. The performance of the recommended model parameters was assessed on 2 sets: (a) the Ln54 benchmark set composed of 54 rare-earth compounds93, and (b) the set of 60 diatomics containing d-block (transition metal) atoms from Ref. 94 (we will refer to this set as Tm60). Due to the challenging SCF orbital optimization landscape and to make the assessment as robust and automated, for each OBS we did assessment on the Ln54 and Tm60 systems for which the exact (four-center) X2C-HF SCF converged in 30 or fewer SCF iterations. The training and assessment was performed with the following OBSs designed for spin-orbit relativity: X2C-SVPall-2c52, X2C-XZVPPall-2c53 (X = T, Q) and all-electron Dyall basis sets dyall-aeYz54, 55, 56 (Y = 2, 3, 4).
The performance of MADF was assessed also against the AutoAux DFBS generator61 as implemented in BSE’s command-line interface41 . For cc-pVXZ, cc-pwCVXZ, and def2- family of OBSs we also compared against the corresponding manually-optimized DFBS cc-pVXZ-RIFIT, cc-pwCVXZ-RIFIT (X = D, T, Q) and def2-SVP-RIFIT and def2-XZVP-RIFIT (X = T, Q) for the G2 set for comparison. All geometries, except for the G2 set, were optimized with the PBE095 and def2-TZVP basis set using the Psi4’s96 geometry optimization module. For elements with atomic number greater than 36, the effective core potentials (ECPs)97 available for def2-TZVP were used for geometry optimization.
All matrix pseudo-inverses and pseudo-inverse square roots were computed with Löwdin’s orthogonalization procedure98.
4 Results
4.1 Training MADF model parameters
The MADF model parameters were tuned to ensure that that the DF errors for the respective training sets did not exceed target accuracies of ±$\pm 20\mu E_{h}$ and ±$\pm 10\mu E_{h}$ per electron for HF and MP2 energies, respectively. This target range for errors was determined by analyzing the DF errors of manually-optimized DFBSs; the majority have larger errors than these targets (see Section 4.2.1).
4.1.1 Training $\tau_{\text{H}}$
The subset of all hydrogen-containing molecules from the TS1 set (Section 3) was used to determine the recommended value of $\tau_{\text{H}}$. While varying $\tau_{\text{H}}$ the remaining MADF parameters were fixed ($\tau_{1}=\tau_{2}=10^{-8},\zeta=1.2$) sufficiently closely to their asymptotic limits ($\tau\to 0$, $\zeta\to 1$) without causing ill-conditioning and excessive costs. Figures 2(a) and 2(b) show the DF errors of HF and MP2 energies, respectively, with $\tau_{\text{H}}$ varying between $10^{-8}$ and $10^{-4}$. The errors are essentially constant with $\tau_{\text{H}}\leq 10^{-6}$, thus $\tau_{\text{H}}=10^{-6}$ is recommended. Since hydrogen has almost negligible relativistic effects, this value will be used henceforth for both relativistic and nonrelativistic calculations.
Figure 2: Variation of DF errors of nonrelativistic (a) HF and (b) MP2 energies of the H-containing subset of the TS1 training set vs the $\tau_{\text{H}}$ model parameter of the MADF generator.
4.1.2 Training $\tau_{1,2}$ and $\zeta$ for nonrelativistic OBS
The variation of DF errors with $\tau_{2}$ was studied next, by keeping the rest of the parameters fixed ($\tau_{1}=10^{-8},\zeta=1.2$) sufficiently closely to their asymptotic limits ($\tau\to 0$, $\zeta\to 1$) without causing ill-conditioning and excessive costs. As Figures 3(a) and 3(b) indicate, the DF errors of HF and MP2 are largely converged at $\tau_{2}=10^{-5}$. Next, $\tau_{2}$ was fixed at $10^{-5}$ and $\tau_{1}$ was varied (Figures 4(a) and 4(b)). The DF errors are sufficiently converged with $\tau_{1}=10^{-6}$. Finally, with $\tau_{1}=10^{-6}$ and $\tau_{2}=10^{-5}$ variation of DF errors with $\zeta$ was studied (Figures 5(a) and 5(b)). For $\zeta>1.4$ the max DF errors increase due to the insufficient spanning of the OBS AO product space by the DFBS AOs, whereas for $\zeta<1.2$ (not shown) the onset of ill-conditioning in DFBS makes the MADF model not numerically stable. $\zeta=1.4$ ensures sufficient convergence of the DF errors without undue numerical problems.
Note that $\zeta=1.4$ is significantly smaller than the exponent ratio 1.8 used by the AutoAux generator ( 1.8)61 and ratio 2 used by PySCF’s generator65 for even-tempered DFBS construction. Our findings confirm that the relatively small exponent ratios are indeed necessary for accurate density fitting. For example, exponent ratios in the $[1.3,1.5]$ range are quite common in the $L\leq 2$ channels of the manually-optimized def2-QZVPP-RIFIT DFBS (see Figure 1). Clearly even-tempered spanning of the AO product space is suboptimal for practical basis sets.
Figure 3: Variation of DF errors of nonrelativistic (a) HF and (b) MP2 energies of the TS1 training set vs the $\tau_{2}$ model parameter of the MADF generator.
Figure 4: Variation of DF errors of nonrelativistic (a) HF and (b) MP2 energies of the TS1 training set vs the $\tau_{1}$ model parameter of the MADF generator.
Figure 5: Variation of DF errors of nonrelativistic (a) HF and (b) MP2 energies of the TS1 training set vs the $\zeta$ model parameter of the MADF generator.
4.1.3 Retraining MADF parameters for relativistic OBS.
Parameters $\tau_{1,2}$ and $\zeta$ were trained again using the relativistic HF and MP2 energies of the TS2 set. $\tau_{2}$ was varied first, with $\tau_{1}=10^{-8}$ and $\zeta=1.2$ (Figures 6(a) and 6(b)); just as in nonrelativistic computations $\tau_{2}=10^{-5}$ ensures sufficient convergence of the DF errors. $\tau_{1}$ was then varied with fixed $\tau_{2}=10^{-5}$ and $\zeta=1.2$ (Figures 7(a) and 7(b)). In contrast to the nonrelativistic case, a much smaller $\tau_{1}$ is needed for sufficient convergence of the DF errors. The variation of relativistic DF errors with $\zeta$ (Figures 8(a) and 8(b)) is similar to the nonrelativistic case, with $\zeta=1.4$ deemed sufficient. Yet again, this is a much smaller ratio than the ratios used by even-tempered DFBS generators.61, 65
The optimal values of the parameters MADF are briefly tabulated in Table 1.
Figure 6: Variation of DF errors of (relativistic) (a) X2C-HF and (b) X2C-MP2 energies of the TS2 training set vs the $\tau_{2}$ model parameter of the MADF generator.
Figure 7: Variation of DF errors of (relativistic) (a) X2C-HF and (b) X2C-MP2 energies of the TS2 training set vs the $\tau_{1}$ model parameter of the MADF generator.
Figure 8: Variation of DF errors of (relativistic) (a) X2C-HF and (b) X2C-MP2 energies of the TS2 training set vs the $\zeta$ model parameter of the MADF generator.
4.2 Assessment of DFBSs generated by the MADF generator.
4.2.1 The G2 set
The DF errors of HF and MP2 energies of the G2 test set produced with DFBS generated by the MADF generator with the recommended model parameters were compared against those with manually-optimized DFBS as well as against those generated by the AutoAux generator. The MADF DFBSs produce smaller DF errors in HF energies than the manually-optimized DFBS (Figure 9(a)), since manually-optimized DFBSs are solely optimized to reduce DF-/RI-MP2 errors. However, the MADF DFBSs also outperform the manually-optimized counterparts for MP2 energies (Figure 9(b)). However, the manually-optimized DFBS are significantly more compact (by as much as a factor of 2) than the MADF counterparts, although the gap decreases significantly for quadruple-zeta basis sets (Figure 9(c)). The comparison to the AutoAux generator is more interesting. Despite fewer parameters, MADF DFBSs produce similar DF errors in HF and MP2 energies. This is especially striking since the MADF DFBS are more compact than the AutoAux counterparts, sometimes by a significant margin (such as for def2-TZVP).
Figure 9: Comparison of DF errors of (a) HF and (b) MP2 energies of the G2 test set obtained with manually-optimized, AutoAux, and MADF DFBSs. (c) Comparison of DFBS sizes, relative to that of the corresponding OBS.
4.2.2 Ln54 and Tm60 sets
The value of DFBS generators is especially pronounced for heavier elements where manually-optimized DFBS are simply not available. Figures 10 and 11 illustrate the relative performance of MADF and AutoAux DFBSs for relativistic computations on molecules with d-block and f-block elements, respectively. It is evident from Figures 10(a) and 11(a) that the DF errors of the X2C-HF energies are smaller with MADF DFBSs than with AutoAux DFBSs, especially for the TURBOMOLE sets. For the MP2 energies (Figures 10(b) and 11(b)) MADF DFBSs are also more accurate, especially for the Dyall basis sets. Although for the TURBOMOLE OBSs MADF DFBS are slightly larger than the AutoAux counterparts, the smaller errors (especially for the triple-zeta OBS) make the increase in the basis set palatable. For the Dyall basis sets MADF generated significantly smaller DFBSs than AutoAux, while producing comparable DF errors of HF energies and much smaller DF errors of MP2 energies.
Figure 10: Comparison of DF errors of (a) X2C-HF and (b) X2C-MP2 energies of the Tm60 test set obtained with AutoAux and MADF DFBSs. (c) Comparison of DFBS sizes, relative to that of the corresponding OBS.
Figure 11: Comparison of DF errors of (a) X2C-HF and (b) X2C-MP2 energies of the Ln54 test set obtained with AutoAux and MADF DFBSs. (c) Comparison of DFBS sizes, relative to that of the corresponding OBS.
5 Summary
This work introduced the MADF algorithm for generating density-fitting basis set composed of primitive solid-harmonic Gaussian AOs and suited for nonrelativistic and relativistic (all-electron) computations with mean-field and correlated electronic structure models. Not using even-tempered AO sets for spanning the AO product space, as in other comparable generators,61, 65 produces compact and accurate candidate pools of DFBS AOs while sufficiently reducing the numerical redundancy. The subsequent pruning of the candidate DFBS pool based on the 2-body energy as the importance metric allows to naturally reduce the relevant range of angular momentum channels and exponent spans with minimal use of heuristics. Generation of primitive (rather than contracted) DF AOs allows to keep the integral evaluation costs optimally low. Using only 4 model parameters MADF generates basis sets that match or exceed the accuracy-to-size ratio of the state-of-the-art DFBS generators; the comparison is particularly favorable with OBSs suited for relativistic all-electron computations. Computations utilizing MADF DFBS, of course, can also benefit from system-specific DFBS compressions.67, 99
Although picture change and other relativistic effects on the 2-particle interaction were not considered here, it will be interesting to examine the impact of such effects on the optimal MADF model parameters in the future; it is known that there are nontrivial differences in the requirements on the DFBS that stem from relativistic contributions to the effective 2-particle interactions.13 Even in the nonrelativistic framework much more testing of the approach will be clearly needed. Most important is to test the performance in the context of higher-order correlated methods (especially coupled cluster), for DF approximation of non-Coulomb operators, for properties other than energy, and for other basis set families (e.g., multiply-augmented correlation-consistent sets, ANO, etc.). Early experiences with MADF DFBSs appear promising. For example, we recently utilized the MADF DFBS generator for explicitly-correlated (F12) coupled-cluster computations with aug-cc-pV7Z OBS (for which no manually-optimized DFBS is available) where we found that the DF errors with the MADF DFBS are significantly smaller than the use of the manually-optimized aug-cc-pV6Z-RIFIT basis. These and other experiences with MADF DFBSs will be reported elsewhere.
{suppinfo}
MADF density-fitting basis sets generated in this work in the Gaussian 94 basis set format; molecular geometries of all species.
{acknowledgement}
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, and Office of Basic Energy Sciences, Scientific Discovery through the Advanced Computing (SciDAC) program under Award DE-SC0022263. This project used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract DE-AC02-05CH11231 using NERSC award ERCAP-0024336. The authors acknowledge Advanced Research Computing at Virginia Tech (https://arc.vt.edu/) for providing computational resources and technical support that have contributed to the results reported within this paper. The development of the Libint software library is supported by the Office of Advanced Cyberinfrastructure, National Science Foundation (Award OAC-2103738).
Appendix A Superposition of atomic densities (SOAD)
Unlike the SAD method,85 with computes the Fock matrix from the atomic (optionally, spin-free) ensemble density matrix computed in OBS self-consistently, SOAD Fock matrix is computed from fixed spin-free ensemble occupancies ${\bf n}_{\mathrm{SOAD}}$ of the minimal basis set (MBS) AOs, without self-consistency. ${\bf n}_{\mathrm{SOAD}}$ for a single atom in an orthonormal MBS is obtained by smearing each subshell’s electrons evenly among all of its orbitals; e.g., for a carbon atom occupancies of 1s, 2s, 2px, 2py, and 2pz MBS AOs are 2, 2, 2/3, 2/3, and 2/3, respectively. This recipe has been referred to as average-of-configuration (AOC)100 and canonical ensemble101. The number of electrons in each subshell is defined by the electron configuration of that atom’s neutral ground state taken from NIST Atomic Spectra Database102. For a molecule ${\bf n}_{\mathrm{SOAD}}$ is obtained by concatenation of the atomic occupancies, without taking into account the effects of nonunit molecular MBS metric. In this work the MINI basis103, 104 was used as MBS for $Z\leq 54$, and the ANO-RCC-MB basis105 for $54<Z\leq 96$.
${\bf n}_{\mathrm{SOAD}}$ is used to construct SOAD Fock matrix in OBS AO basis, $\mathbf{F_{\mathrm{SOAD}}}$, as follows:
$(\mathbf{F}_{\mathrm{SOAD}})^{\mu}_{\nu}=h^{\mu}_{\nu}+\sum_{m}({\bf n}_{\mathrm{SOAD}})_{m}[2(\mu\nu|mm)-(\mu m|\nu m)],$ (25)
where $h^{\mu}_{\nu}$ are the matrix elements of the nonrelativistic 1-body (core) Hamiltonian, $m$ runs over MBS AOs and $\mu,\nu$ over OBS AOs.

### v2

@@
 The algorithm is designed to (a) saturate the OBS product space by a large regularized set of primitive solid-harmonic Gaussian shells with nonuniform distribution of exponents followed by (b) pruning of the shells according to their contributions to the 2-body energy of a correlated atomic ensemble.
-Building the DFBS generator model almost exclusively on mathematical and physical principles allows one to limit the number of parameters that control the density fitting error to four, with a single set of parameters sufficient for computations with all basis cardinal numbers, with and without correlation of core electrons, with and without scalar and spin-dependent relativistic effects, spanning almost all of the Periodic Table.
+Building the DFBS generator model almost exclusively on mathematical and physical principles allows one to limit the number of parameters that control the density fitting error to three, with a single set of parameters sufficient for computations with all basis cardinal numbers, with and without correlation of core electrons, with and without scalar and spin-dependent relativistic effects, spanning almost all of the Periodic Table.
 Performance assessment included basis sets up to quadruple-zeta quality from several major basis set families, using molecules composed of main-group, d-block, and f-block elements.
-The resulting DF errors in Hartree-Fock and second-order MP2 energies (with relativistic all-electron treatments, when appropriate) were on the order of $20$ and $10$ $\mu E_{h}$ per electron, respectively.
+The resulting DF errors in Hartree-Fock and second-order MP2 energies (with relativistic all-electron treatments, when appropriate) were on the order of $20$ and 1010 $\mu E_{h}$ per electron, respectively.
 1 Introduction
@@
 Most DBFS generators are designed only for DF in the context of evaluating the Coulomb potential of a charge density.
-These include the AutoABS method by Yang et al.57, long-range corrected auxiliary basis sets by Hellmann et al.58, even tempered auxiliary basis sets with shared exponents by Daz-Tinoco et al.59 and a similar approach has also recently been reported for relativistic OBSs60.
+These include the AutoABS method by Yang et al.57, atomic Cholesky decomposition (aCD) and atomic compact Cholesky decomposition (acCD)64, 65, long-range corrected auxiliary basis sets by Hellmann et al.58, even tempered auxiliary basis sets with shared exponents by Daz-Tinoco et al.59 and a similar approach has also recently been reported for relativistic OBSs60.
 DFBS generators designed for mean-field and correlated electronic structure simulation include the AutoAux procedure by Stoychev et al.,61 methods proposed by Lehtola62, 63, and PySCF built-in generator of even-tempered DFBSs.
-The AutoAux method, available in the ORCA software package64 and in the command-line interface of Basis Set Exchange (BSE)41, produces a primitive even-tempered DFBS that can be applied universally (in mean-field and correlated simulations).
+The AutoAux method, available in the ORCA software package66 and in the command-line interface of Basis Set Exchange (BSE)41, produces a primitive even-tempered DFBS that can be applied universally (in mean-field and correlated simulations).
 The AutoAux generator uses even-tempered sequences of Gaussian primitive AOs to span the OBS-deduced range of exponents for each angular momentum block.
@@
 AutoAux as a result is quite complex, defined with more than a dozen model parameters in total.
-PySCF’s simple built-in generator65 produces even-tempered primitive DFBS and is very similar in spirit to AutoAux.
+PySCF’s simple built-in generator67 produces even-tempered primitive DFBS and is very similar in spirit to AutoAux.
 It has fewer model parameters (e.g., even-tempering ratio is not graded with angular momenta) and claims to offer similar accuracy to AutoAux, although no extensive comparison exists.
@@
 Although the resulting DFBSs are nearly exact, they are extremely large, lead to high condition numbers in molecules, and contain primitive Gaussians with twice the angular momentum of the highest angular momentum ($L_{\text{OBS}}$) of functions in the parent OBS.
-Lehtola addressed the latter issue in his recent work66 by extending the algorithm to prune high-angular momentum functions out of the DFBS with a tunable parameter.
-The algorithm then uses Kállay’s67 contraction method with singular value decomposition (SVD) to contract the DFBS, further reducing its size.
-An attractive alternative to density fitting that avoids these issues and provides some of its advantages is the (pivoted) Cholesky decomposition (CD).68 Any positive-definite 2-particle interaction represented in an arbitrary OBS can be decomposed with accuracy controlled by a single threshold.
-CD has been used for electronic structure simulations in both relativistic69, 70, 71 and nonrelativistic72, 73, 74 contexts.
-Although the accuracy of CD can be robustly pushed beyond the reach of conventional density fitting, the cost of CD is substantial.
-Computing analytic derivatives of the Cholesky vectors is also fundamentally challenging.
-The 1-center CD (1C-CD),75 later renamed atomic CD (aCD),76 works around the problem of analytic derivatives of CD by marrying CD with density fitting.
-Namely, global CD is avoided and CD is only used to construct a contracted DFBS for each atom suitable for accurate representation of a given operator in a given OBS.
-Thus the aCD approach can be viewed as a black-box 1-parameter DFBS generator.
-The aCD approach also produces contracted DFBSs which have mixed angular momentum functions and mixed representations (cartesian or spherical).
-A solution to this issue was proposed by Hellmann et al.77, combining aCD and Lehtola’s method62, however, it suffers from the same issues as Lehtola’s original method.62 Several challenges in designing efficient CD algorithms in the context of quantum chemistry have also been mentioned.78
-To navigate the severe gaps in the coverage by manually-constructed DFBS of existing work and the significant shortcomings of the existing DFBS generators we attempted to design a new DFBS generator that produces (a) primitive Gaussian DFBSs, (b) of comparable size and accuracy as the manually-optimized DFBS, (c) usable in the context of mean-field and correlated simulation, and (d) has as few adjustable parameters as possible to ensure maximum universality.
+Hellmann et al.68 extended Lehtola’s method62 to include exact spherical functions with mixed angular momentum functions instead of creating multiple angular momentum channels; however, it does not address the large number of primitives in the generated DFBS.
+Lehtola addressed the latter issue in his recent work69 by extending the algorithm to prune high-angular momentum functions out of the DFBS with a tunable parameter.
+The algorithm then uses Kállay’s70 contraction method with singular value decomposition (SVD) to contract the DFBS, further reducing its size.
+An attractive alternative to density fitting that avoids these issues and provides some of its advantages is the (pivoted) Cholesky decomposition (CD).71 Any positive-definite 2-particle interaction represented in an arbitrary OBS can be decomposed with accuracy controlled by a single threshold.
+CD has been used for electronic structure simulations in both relativistic72, 73, 74 and nonrelativistic75, 76, 77 contexts and close connections between CD and DF have been demonstrated78.
+Although the accuracy of CD can be robustly pushed beyond the reach of conventional density fitting, the cost of molecular CD can be substantially higher than that of DF-based approaches and CD is not as universally applicable black-box technology as desired.79 The use of CD for construction of atomic density fitting basis sets, under the name of atomic CD (aCD), was pioneered by Aquilante et al.64 The improved version of aCD, termed atomic compact CD (acCD),80 was proposed later that greatly reduced the number of primitive and contracted DF AOs.
+Both the aCD and acCD approaches can be viewed as black-box 1-parameter DFBS generators and are also unbiased towards mean-field and correlated methods, albeit they produce deeply contracted DFBS AOs.
+To navigate the severe gaps in the coverage by existing manually-constructed DFBS and the notable shortcomings of the existing DFBS generators we attempted to design a new DFBS generator that produces (a) primitive Gaussian DFBSs, (b) of comparable size and accuracy as the manually-optimized DFBS, (c) usable in the context of mean-field and correlated simulation, and (d) has as few adjustable parameters as possible to ensure maximum universality.
 By combining the existing ideas for spanning the product space of OBS AOs with new ideas for regularizing and pruning the exact DFBS using atomic correlated ensemble RDMs we arrived at a DFBS generator dubbed MADF.
@@
 Section 3 describes the technical and details of the parameter training regiment.
-In Section 4 we document the performance of the DFBSs generated by our method against the manually-optimized DFBSs and the DFBSs generated by the AutoAux algorithm.
+In Section 4 we document the performance of the DFBSs generated by our method against the manually-optimized DFBSs and the DFBSs generated by the AutoAux algorithm and acCD.
 2 Formalism
@@
 DF approximates product of AOs (“density”) $\phi_{\mu}({\bf r})\phi_{\nu}({\bf r})\equiv({\bf r}|\mu\nu)$ as a linear combination of DFBS AOs $\phi_{X}({\bf r})$
-$|\mu\nu)\approx\sum_{X}C_{\mu\nu}^{X}|X)\equiv|\widetilde{\mu\nu}).$ (1)
-The fitting coefficients $C_{\mu\nu}^{X}$ are obtained by minimizing a norm of the density error $|\delta_{\mu\nu})\equiv|\mu\nu)-|\widetilde{\mu\nu})$.
-To minimize the error in the diagonal elements of the matrix representation of a positive operator $\hat{O}$, namely in $(\mu\nu|\hat{O}|\mu\nu)$, the optimal choice of the error norm is the expectation value of operator $\hat{O}$ itself:
-$||f||_{\hat{O}}\equiv(f|\hat{O}|f),$ (2)
-This is known as robust density fitting.79, 80 Least-squares minimization of $||\delta_{\mu\nu}||_{\hat{O}}$ is equivalent to solving the following equation
+[…]
+[…]
+[…]
+Such choice makes the error in $(\mu\nu|\hat{O}|\rho\sigma)$ quadratic in the density fitting errors and is a special case of robust density fitting;81, 82 if the fitting norm were defined with operator $\hat{W}\neq\hat{O}$ the explicit robust fitting approach would be necessary and variational extensions83 could be introduced for convenience.
+Least-squares minimization of $\delta_{\mu\nu}||_{\hat{O}}$ is equivalent to solving the following equation
 $\hat{O}|\mu\nu)=\sum_{Y}\hat{O}|Y)C^{Y}_{\mu\nu}.$ (3)
@@
 the robust fitting minimizes the $L^{2}$ norm of the fitting error in the electric field generated by the AO product.
-DFBS generators usually57, 58, 59, 60, 61, 62, 63, 64 e.g., AutoAux, aCD work in a single-atom context, i.e., DFBS is constructed using information about products on OBS AOs on a single atom.
+DFBS generators usually57, 58, 59, 60, 61, 62, 63, 66 work in a single-atom context, i.e., DFBS is constructed using information about products on OBS AOs on a single atom.
 It is reasonable to question whether this is sufficient.
 Assuming a system composed of a single atom type, global density fitting in a system of $n$ atoms will be able to fit every such 1-center product of OBS AOs using $n$ times as many DFBS AOs as in a single atom.
-Indeed, for long-range $\hat{O}$ the fitting of even strongly-localized AO products generally involves DFBS AOs from far away.81 Thus, 1-center products in a molecule can borrow DFBS AOs from other centers to make the fitting more accurate than in a single atom.
+Indeed, for long-range $\hat{O}$ the fitting of even strongly-localized AO products generally involves DFBS AOs from far away.84 Thus, 1-center products in a molecule can borrow DFBS AOs from other centers to make the fitting more accurate than in a single atom.
 This suggests that a single atom poses a more challenging DF setting than a molecule, which supports the use of a single-atom setting for the generation of DFBS.
@@
 Furthermore, we will limit ourselves to the case of OBS and DFBS composed of solid-harmonic Gaussian (SHG) AOs.
-Unnormalized primitive SHG with exponent $\alpha$ is defined as:82, 83
+Unnormalized primitive SHG with exponent $\alpha$ is defined as:85, 86
 $\chi_{\alpha,l,m}({\bf r})=$ $r^{l}\exp(-\alpha r)y_{lm}(\theta,\phi)$ (6)
@@
 Clearly, the contribution to the product from the Clebsch-Gordan channel $L<l_{\mu}+l_{\nu}$ is an SHG multiplied by $r^{l_{\mu}+l_{\nu}-L}$.
-Lehtola proposed to account for this effect by approximating each such contribution by a single SHG of angular momentum $L$ with the following effective exponent (see Appendix II of Ref.
+Following Stoychev61, Lehtola proposed to account for this effect by approximating each such contribution by a single SHG of angular momentum $L$ with the following effective exponent (see Appendix II of Ref.
 62):
@@
 The set $\mathcal{C}$ can be used as DFBS for nearly exact density fitting in an atom (a) for any contraction of the generating set of primitive OBS AOs and (b) for any operator.
-However, even in a single atom such DFBS is highly linearly dependent and overcomplete, 84, 62 therefore, it is unusable with finite precision arithmetic.
+However, even in a single atom such DFBS is highly linearly dependent and overcomplete, 87, 62 therefore, it is unusable with finite precision arithmetic.
 Moreover, such a basis would be extremely large, defeating the purpose of the DF approximation.
@@
 Figure 1(a) shows distribution of DFBS exponents for nonrelativistic OBS def2-QZVPP and Figure 1(b) for relativistic OBS x2c-QZVPPall-2c.
-The plot depict Lehtola’s complete set (shown in purple), Lehtola’s pCD-regularized set (shown in red) with pCD threshold $10^{-7}$ as recommended in Ref.62, even-tempered exponents produced by AutoAux (shown in green) and set of exponents produced by the MADF algorithm (shown in orange).
+The plot depict Lehtola’s complete set (shown in purple), Lehtola’s pCD-regularized set (shown in red) obtained using ERKALE88 with pCD threshold $10^{-7}$ as recommended in Ref.62, even-tempered exponents produced by AutoAux (shown in green) and set of exponents produced by the MADF algorithm (shown in orange).
 Distribution of exponents of manually-optimized DFBS def2-QZVPP-RIFIT is also shown (in blue) in Figure 1(a).
@@
 Given a primitive SHG set $\mathcal{C}$ and a target exponent ratio threshold $\zeta$, each subset $\mathcal{C}_{L}$ containing primitives of angular momenta $L$ is regularized by repeatedly replacing the pair of primitive SHG shells with the smallest exponent ratio $\alpha_{1}/\alpha_{2}$ by a single primitive SHG shell with the geometric mean exponent $\alpha=\sqrt{\alpha_{1}\alpha_{2}}$.
+If multiple exponent pairs have a “soft” tie for the smallest exponent ratio, the pair with the smallest orbital exponent wins.
 Regularization stops if no pairs of primitives have an exponent ratio less than $\zeta$.
@@
 1:function RegularizedCompleteDFBS($\mathcal{P}$, $\zeta$) $\triangleright$ $\mathcal{P}$: primitive OBS shells; $\zeta\geq 1$
-2:     $\mathcal{C}\leftarrow\textsc{CompleteDFBS}(\mathcal{P})$triangleright Complete set of DF SHG shells with exponents generated by Eq. 8
-3:     $L_{\text{max}}\leftarrow\textsc{MaxAM}(\mathcal{c})$triangleright Maximum angular momentum of shells in $\mathcal{c}$
-4:     ${}\mathcal{R}\leftarrow\{\}$triangleright Initialize the regularized complete DFBS
+2:  $\mathcal{C}\leftarrow\textsc{CompleteDFBS}(\mathcal{P})$ $\triangleright$ Complete set of DF SHG shells with exponents generated by Eq. 8
+3:  $L_{\text{max}}\leftarrow\textsc{MaxAM}(\mathcal{c})$ $\triangleright$ Maximum angular momentum of shells in $\mathcal{c}$
+4:  $\mathcal{R}\leftarrow\{\}$ $\triangleright$ Initialize the regularized complete DFBS
 5:  for $L\in[0,L_{\text{max}}]$ do
-6:         $\mathcal{c}\leftarrow\textsc{SelectByAM}(\mathcal{C},L)$triangleright Shells in $\mathcal{C}$ with angular momentum $L$
-7:         if Size($\mathcal{c}$)$\geq 2$ then
-8:              $\zeta_{\text{min}}\leftarrow\infty$triangleright smallest exponent ratio
+6:   $\mathcal{c}\leftarrow\textsc{SelectByAM}(\mathcal{C},L)$ $\triangleright$ Shells in $\mathcal{C}$ with angular momentum $L$
+[…]
+8:     $\zeta_{\text{min}}\leftarrow\textsc{MaxRatio}(\mathcal{c})$ $\triangleright$ smallest exponent ratio
 9:     $I_{\text{min}}\leftarrow 0$
 10:     while $\zeta_{\text{min}}<\zeta$ do
-11:                  $\mathcal{c}\leftarrow\textsc{SortByExponent}(\mathcal{c})$triangleright Sort in order of decreasing exponents
-12:                  for $i$ in $20\dots\textsc{Size}(\mathcal{c})-$ 2 do
-13:                       $\zeta_{i}\leftarrow\textsc{Exponent}(\mathcal{c}\left[i\right])/\textsc{Exponent}(\mathcal{c}\left[i+1\right])$
-14:                       if $\zeta_{i}<\zeta_{\text{min}}$ then
-15:                           $\zeta_{\text{min}}\leftarrow\zeta_{i}$
-16:                           $I_{\text{min}}\leftarrow i$triangleright Tracks shell pair with smallest exponent ratio
-17:                       else
-18:                           continue
-19:                       end if
-20:                  end for
-21:                  $\alpha_{I}\leftarrow\textsc{Exponent}(\mathcal{c}\left[I_{\text{min}}\right])$
-22:                  $\alpha_{I+1}\leftarrow\textsc{Exponent}(\mathcal{c}\left[I_{\text{min}}+1\right])$
-23:                  $c\leftarrow\textsc{SHGShell}(\sqrt{\alpha_{I}\alpha_{I+1}},L)$triangleright Pair of shells with closest exponents fused to single shell
-[…]
-25:              end while
-26:         end if
-27:         $\mathcal{R}\leftarrow\textsc{Concat}(\mathcal{R},\mathcal{c})$
-28:     end for
-29:     return $\mathcal{R}$
-30:end function
+11:      $\mathcal{c}\leftarrow\textsc{SortByExponent}(\mathcal{c})$ $\triangleright$ Sort in order of decreasing exponents
+12:      $\mathcal{p}\leftarrow\textsc{Neighbors}(\mathcal{c})$ $\triangleright$ List of {Exponent($\mathcal{c}[i]$), Exponent($\mathcal{c}[i+1]$), ii} tuples
+13:      $\mathcal{p}\leftarrow\textsc{SortByExponentRatio}(\mathcal{p})$ $\triangleright$ Sort in order of increasing exponent ratios
+14:      $\zeta_{\text{min}}\leftarrow\mathcal{p}[0,0]/\mathcal{p}[0,1]$
+15:      $\mathcal{p}\leftarrow\textsc{MinRatioTies}(\mathcal{p},\zeta_{\text{min}})$ $\triangleright$ Select pairs with exponent ratios in $\zeta_{\text{min}},\zeta_{\text{min}}+10^{-8}]$
+16:      $\{\alpha_{I},\alpha_{I+1},I_{\text{min}}\}\leftarrow\textsc{MinElement}(\mathcal{p})$ $\triangleright$ Find the pair with the smallest exponents
+17:      $c\leftarrow\textsc{SHGShell}(\sqrt{\alpha_{I}\alpha_{I+1}},L)$ $\triangleright$ Pair of shells with closest exponents fused to single shell
+18:      $\mathcal{c}\leftarrow$ Concat($\mathcal{c}\left[0:I_{\text{min}}-1\right],c,\mathcal{c}\left[I_{\text{min}}+2:-1\right]$)
+19:     end while
+20:   end if
+21:   $\mathcal{R}\leftarrow\textsc{Concat}(\mathcal{R},\mathcal{c})$
+22:  end for
+23:  $\mathcal{R}\leftarrow\textsc{CholeskyScreen}(\mathcal{R},10^{-8})$ $\triangleright$ See Ref.62, only drops strongly linearly dependent primitives.
+24:  return $\mathcal{R}$
+25:end function
 2.3 Estimating energy contribution of DFBS primitives
@@
 In ad hoc DFBS generators like AutoAux, the angular momenta of DFBS are restricted heuristically, i.e.
-$L_{\text{max,DF}}=\text{max}(2L_{\text{occ}},L_{\text{OBS}}+L_{\text{inc}})$, where $L_{\text{inc}}$ is fixed to 1 or 2 depending on $Z$61 To avoid the heuristics as much as possible, we attempted to design a first-principles energy-based model for further pruning of DFBS.
+$L_{\text{max,DF}}=\text{max}(2L_{\text{occ}},L_{\text{OBS}}+L_{\text{inc}})$, where $L_{\text{inc}}$ is fixed to 1 or 2 depending on ZZ.61 To avoid the heuristics as much as possible, we attempted to design a first-principles energy-based model for further pruning of DFBS.
 Consider the expectation value of a 2-body Fock-space Hamiltonian:
@@
 Thus in practice we constructed an estimator that requires a model for orbital occupancies $n_{p}\equiv\gamma_{p}^{p}$ only, rather than the full 2-RDM.
-To model the 2-RDM in terms of orbital occupancies, we use the Cauchy-Schwarz inequalities for the 2-RDM,82 which for positive $\gamma^{pq}_{pq}$ becomes simply:
+To model the 2-RDM in terms of orbital occupancies, we use the Cauchy-Schwarz inequalities for the 2-RDM,85 which for positive $\gamma^{pq}_{pq}$ becomes simply:
 $\gamma^{pq}_{pq}\leq\sqrt{n_{p}n_{q}}.$ (13)
@@
 This expression is independent of whether the single-particle states are spin-free or spin-orbital. Since robust DF guarantees that the error in $(pq|qp)=(pq|pq)$ is positive:
-[…]
+$(pq|pq)-(\widetilde{pq}|\widetilde{pq})\overset{\text{robust~DF}}{\geq}0.$ (15)
 then the robust DF approximation to Eq. 14,
@@
 For an atom with ground-state configuration including occupied subshells of angular momenta $\leq L_{\text{occ}}$ DF channels $L\leq 2L_{\text{occ}}$ are pruned with $\tau=\tau_{1}$, the rest are pruned with $\tau=\tau_{2}$.
-Using the same values of $\tau$ for $Z=1$ (hydrogen) as for other $Z$ produced larger errors than desired due to the overpruning of the higher $L$ DF channels; thus, separate threshold $\tau_{\text{H}}$ is used for pruning DF channels of all angular momenta for H.
-The values of $\tau_{1}$, $\tau_{2}$, and $\tau_{\text{H}}$ will be determined heuristically in Section 4.1.
+For atoms with $L_{\text{occ}}=0$, namely H, He, Li and Be, to avoid overpruning of the higher $L$ DF channels, we consider $L_{\text{occ}}$ to be 1.
+The values of $\tau_{1}$ and $\tau_{2}$will be determined heuristically in Section 4.1.
 2.4 Correlated model of orbital occupancies
@@
 Such definition for the OBS Fock matrix is utilized in the Libint library and the MPQC program under the name “superposition of atomic densities” (SOAD) for generating guess Fock matrices for SCF.
-SOAD can be viewed as a simplified version of the popular SAD method85, with differences described for completeness in appendix A.
+SOAD can be viewed as a simplified version of the popular SAD method89, with differences described for completeness in Appendix A.
 Diagonalization of SOAD ensemble Fock matrix expressed in OBS AO (see Appendix A),
@@
 produces energies and OBS AO coefficients of non-self-consistent ensemble orbitals.
-The orbitals are assigned spin-free occupancies $0\leq{\bf n}^{(0)}\leq 2$ using the same spherically-averaged ensemble definition used to construct the SOAD ensemble density and $\mathbf{F}_{\mathrm{SOAD}}$.
+In this work, we have used the spin-free exact two-component 1-body (core) Hamiltonian (sf-1eX2C)90, 91, 92 to construct the $\mathbf{F}_{\mathrm{SOAD}}$ matrix to account for scalar relativistic effects.
+In general the model orbital energies of the atomic subshells obtained from Eq.
+20 do not match the canonical order of Aufbau principle (Madelung rule) or even the order produced by self-consistent field in the given OBS.
+Thus the uncorrelated occupation numbers for the OBS ensemble orbitals ${\bf n}^{(0)}$ are obtained by mapping the spherically-averaged MBS ensemble occupancies ${\bf n}_{\mathrm{SOAD}}$ (Appendix A) via:
+${\bf n}^{(0)}=\mathrm{id}\left(\mathbf{C}^{\dagger}\mathbf{S}_{\mathrm{OBS,MBS}}\right){\bf n}_{\mathrm{SOAD}},$ (21)
+where $\mathbf{S}_{\mathrm{OBS,MBS}}$ is the overlap of OBS and MBS AOs, and $\mathrm{id}$ maps the real-valued matrix argument to the nearest identity matrix.
 Columns in $\bf C$} corresponding to zero and nonzero ensemble occupancies are deemed unoccupied ($a,b$) and occupied ($i,j$), respectively.
 These orbitals and the corresponding orbital energies are used to compute opposite-spin ensemble Møller-Plesset first-order amplitudes:
-$t_{ab}^{ij}=\frac{\sqrt{n_{i}^{(0)}n_{j}^{(0)}}}{2}\frac{(ia|jb)}{\varepsilon_{i}+\varepsilon_{j}-\varepsilon_{a}-\varepsilon_{b}}.$ (21)
-The corresponding second-order correlation contributions ${\bf n}^{(2)}$ to orbital occupancies are obtained straghtforwardly:
+$t_{ab}^{ij}=-\frac{\sqrt{n_{i}^{(0)}n_{j}^{(0)}}}{2}\frac{(ia|jb)}{\varepsilon_{a}+\varepsilon_{b}-\varepsilon_{i}-\varepsilon_{j}}(1-e^{-\kappa(\varepsilon_{a}+\varepsilon_{b}-\varepsilon_{i}-\varepsilon_{j})})^{2}.$ (22)
+The last factor on the right-hand side is the $\kappa$-regularizer used by Lee and Head-Gordon93 in the context of orbital-optimized MP2 method.
+Whereas in Ref.
+93 heuristic tuning suggested $\kappa=1.45\,E_{\mathrm{h}}^{-1}$ as best at avoiding the singularities and accounting for the inaccuracy of the MP1 amplitude model.
+Here we set $\kappa=3\,E_{\mathrm{h}}^{-1}$ as the purpose of the regularizer here is to only avoid potential singularities due to possible degeneracies between occupied and unoccupied subshells; the inadequacy of the MP1 model for describing correlations of electrons in the partially occupied subshells is already partially addressed by avoiding internal excitations within such subshells.
+The corresponding second-order correlation contributions ${\bf n}^{(2)}$ to orbital occupancies are obtained straightforwardly:
 $n_{i}^{(2)}=$ $\sum_{jab}-2t^{ij}_{ab}t_{ij}^{ab},$ (23)
@@
 For a given (input) OBS consisting of contracted SHG AOs on an atom with atomic number $Z$ MADF (1) generates a “complete” set of primitive DF SHG AOs necessary to represent all products of uncontracted OBS AOs using the method of Lehtola62, (2) regularizes it by fusing pairs of shells with overly close exponents as described in Section 2.2, and (3) prunes out DF AOs not important for description of correlated 2-body energy using the method described in Section 2.3.
-Complete description of the algorithm is given in Algorithm 2.
-Its 4 model parameters (with 5 unique values) summarized in Table 1.
+The complete description of the algorithm is given in Algorithm 2.
+Its 3 model parameters (with 4 unique values) are summarized in Table 1.
 Parameter Description Values
 $\zeta$ Exponent ratio threshold for Algorithm 1 1.4
-$\tau_{\text{H}}$ 2-body energy threshold for the hydrogen atom $10^{-6}$
-$\tau_{1}$ 2-body energy threshold for $Z>1$ and $L\leq 2l^{occ}$ $10^{-6}$ or $10^{-8}$ [^1] Optimal values of $\tau_{1}$ are different for nonrelativistic and relativistic OBSs, see Sections 4.1.2 and 4.1.3
-$\tau_{2}$ 2-body energy screening threshold for $Z>1$ and $L>2l^{occ}$ $10^{-5}$
+$\tau_{1}$ 2-body energy threshold for $L\leq 2l^{occ}$ $10^{-6}$ (nr), $10^{-7}$ (rel) a
+$\tau_{2}$ 2-body energy screening threshold for $L>2l^{occ}$ $10^{-5}$
 Table 1: Model parameters of the MADF DFBS generator.
+  • a
+    Optimal values of $\tau_{1}$ are different for nonrelativistic and relativistic OBSs, see Sections 4.1.1 and 4.1.2
 Algorithm 2 The MADF algorithm
-1:function MADF($Z,\mathcal{O},\mathcal{M},\tau_{1},\tau_{2},\tau_{\text{H}},\zeta$) $\triangleright$ $Z$: atomic number; $\mathcal{O}$: (contracted) OBS; $\mathcal{M}$: MBS; $\tau_{\{1,2,\text{H}\}}\geq 0$; $\zeta\geq 1$
+1:function MADF($Z,\mathcal{O},\mathcal{M},\tau_{1},\tau_{2},\zeta$) $\triangleright$ $Z$: atomic number; $\mathcal{O}$: (contracted) OBS; $\mathcal{M}$: MBS; $\tau_{\{1,2\}}\geq 0$; $\zeta\geq 1$
 2:  $\mathcal{P}\leftarrow\textsc{Uncontract}(\mathcal{O})$
-3:     $\mathcal{C}\leftarrow\textsc{RegularizedCompleteDFBS}(\mathcal{P},\zeta)$triangleright Algorithm 1
-4:     $\mathbf{F}_{\mathrm{SOAD}}\leftarrow\textsc{SoadFock}(Z,\mathcal{O},\mathcal{M})$triangleright appendix A
-5:     $\mathbf{F}_{\mathrm{SOAD}}\mathbf{C}=\mathbf{SC\epsilon}$triangleright Eq. 20
-6:     ${\bf n}\leftarrow\textsc{CorrelatedOccupancies}(\mathbf{C},\mathcal{M})$triangleright Eqs. 21, 22 and 24
-7:     $L_{\text{occ}}\leftarrow\textsc{MaxAM}(\mathcal{M})$triangleright maximum angular momentum of occupied subshells.
-8:     $L_{\text{max}}\leftarrow\textsc{MaxAM}(\mathcal{R})$triangleright maximum angular momentum of candidate pool of shells
-9:     $\mathcal{D}\leftarrow\{\}$triangleright Initialize the target list of DF shells
+3:  $\mathcal{C}\leftarrow\textsc{RegularizedCompleteDFBS}(\mathcal{P},\zeta)$ $\triangleright$ Algorithm 1
+4:  $\mathbf{F}_{\mathrm{SOAD}}\leftarrow\textsc{SoadFock}(Z,\mathcal{O},\mathcal{M})$ $\triangleright$ Appendix A
+5:  $\mathbf{C}\leftarrow\mathbf{F}_{\mathrm{SOAD}}\mathbf{C}=\mathbf{SC\epsilon}$ $\triangleright$ Eq. 20
+6:  ${\bf n}\leftarrow\textsc{CorrelatedOccupancies}(\mathbf{C},\mathcal{M})$ $\triangleright$ Eqs. 21, 22, 23 and 25
+7:  $L_{\text{occ}}\leftarrow\textsc{Max}(\textsc{MaxAM}(\mathcal{M}),1)$ $\triangleright$ maximum angular momentum of occupied subshells.
+8:  $L_{\text{max}}\leftarrow\textsc{MaxAM}(\mathcal{R})$ $\triangleright$ maximum angular momentum of candidate pool of shells
+9:  $\mathcal{D}\leftarrow\{\}$ $\triangleright$ Initialize the target list of DF shells
 10:  for $L\in[0,L_{\text{max}}]$ do
 11:   $\mathcal{d}\leftarrow\{\}$
-12:         if $Z\neq 1$ then
-13:              if $L\leq 2L_{\text{occ}}$ then
-14:                  $\tau\leftarrow\tau_{1}$
-15:              else
-16:                  $\tau\leftarrow\tau_{2}$
-17:              end if
-18:         else
-19:              $\tau\leftarrow\tau_{\text{H}}$
-20:         end if
-21:         $\mathcal{r}\leftarrow\textsc{SelectByAM}(\mathcal{R},L)$triangleright Shells in $\mathcal{R}$ with angular momentum $L$
-22:         $\{E_{\mathcal{r}},\{\epsilon_{\mathcal{r}}\}\}\leftarrow\textsc{Eval2EdxL}(\mathcal{r})$triangleright Evaluate Eqs. 17 and 18 using $\mathcal{r}$ as DFBS
-23:         if $E_{\mathcal{r}}\geq Z\tau$ then
-24:              $\mathcal{r}\leftarrow\textsc{SortBy}(\mathcal{r},\{|\epsilon_{\mathcal{r}}|\})$triangleright Sort $\mathcal{r}$ in order of decreasing $|{}^{2}\tilde{\bar{\epsilon}}^{\text{dx}}_{X}|$
-25:              $E_{\mathcal{d}}\leftarrow 0,i\leftarrow 0$
-26:              while $E_{\mathcal{r}}-E_{\mathcal{d}}\geq Z\tau$ and $i<\textsc{Size}(r)$ do
-27:                  $\mathcal{d}\leftarrow\textsc{Concat}(\mathcal{d},\mathcal{r}[i])$
-28:                  $\{E_{\mathcal{d}},\_\}\leftarrow\textsc{Eval2EdxL}(\mathcal{d})$triangleright Evaluate Eq. 17 using $\mathcal{d}$ as DFBS
-29:                  $i\leftarrow i+1$
-30:              end while
-31:              $\mathcal{D}\leftarrow\textsc{Concat}(\mathcal{D},\mathcal{d})$
-32:         end if
-33:     end for
-34:     return $\mathcal{D}$
-35:end function
+12:   if $L\leq 2L_{\text{occ}}$ then
+13:     $\tau\leftarrow\tau_{1}$
+14:   else
+15:     $\tau\leftarrow\tau_{2}$
+16:   end if
+17:   $\mathcal{r}\leftarrow\textsc{SelectByAM}(\mathcal{R},L)$ $\triangleright$ Shells in $\mathcal{R}$ with angular momentum $L$
+18:   $\{E_{\mathcal{r}},\{\epsilon_{\mathcal{r}}\}\}\leftarrow\textsc{Eval2EdxL}(\mathcal{r})$ $\triangleright$ Evaluate Eqs. 17 and 18 using $\mathcal{r}$ as DFBS
+19:   if $E_{\mathcal{r}}\geq Z\tau$ then
+20:     $\mathcal{r}\leftarrow\textsc{SortBy}(\mathcal{r},\{|\epsilon_{\mathcal{r}}|\})$ $\triangleright$ Sort $\mathcal{r}$ in order of decreasing $|{}^{2}\tilde{\bar{\epsilon}}^{\text{dx}}_{X}|$
+21:     $E_{\mathcal{d}}\leftarrow 0,i\leftarrow 0$
+22:     while $E_{\mathcal{r}}-E_{\mathcal{d}}\geq Z\tau$ and $i<\textsc{Size}(r)$ do
+23:      $\mathcal{d}\leftarrow\textsc{Concat}(\mathcal{d},\mathcal{r}[i])$
+24:      $\{E_{\mathcal{d}},\_\}\leftarrow\textsc{Eval2EdxL}(\mathcal{d})$ $\triangleright$ Evaluate Eq. 17 using $\mathcal{d}$ as DFBS
+25:      $i\leftarrow i+1$
+26:     end while
+27:     $\mathcal{D}\leftarrow\textsc{Concat}(\mathcal{D},\mathcal{d})$
+28:   end if
+29:  end for
+30:  return $\mathcal{D}$
+31:end function
 3 Technical details
-The MADF algorithm is implemented in a developmental version of MPQC86; generation of base DFBS by Lehtola’s algorithm62 was implemented in Libint.87, 88 For optimization of MADF parameters and its benchmark testing, we employed the nonrelativistic and relativistic Hartree-Fock (HF) and second-order Møller-Plesset (MP2) methods.
+The MADF algorithm is implemented in a developmental version of MPQC94; generation of base DFBS by Lehtola’s algorithm62 was implemented in Libint.95, 96 For optimization of MADF parameters and its benchmark testing, we employed the nonrelativistic and relativistic Hartree-Fock (HF) and second-order Møller-Plesset (MP2) methods.
 DF errors were computed relative to the HF and MP2 energies obtained with exact (four-center) two-electron matrix elements throughout.
-The exact two-component decoupled one-body Hamiltonian89, 90, 91 (1eX2C, or simply X2C) with spin-orbit coupling was used for all relativistic computations without empirical scaling corrections.
+The exact two-component decoupled one-body Hamiltonian90, 91, 92 (1eX2C, or simply X2C) with spin-orbit coupling was used for all relativistic computations without empirical scaling corrections.
 The core orbitals were frozen in nonrelativistic MP2 computations with cc-pVXZ and def2- OBSs.
 The MADF model parameters for use with nonrelativistic OBS were determined on training set TS1 composed of 38 closed-shell systems with light elements: AlF3, Ar, BF3, BH3, C2H2, C2H4, C2H6, CH3OH, CH4, Cl2, CO2, COH2, CS2, F2, H2O, HCl, HF, HNO2, N2, cis-N2H2, trans-N2H2, Ne, NH3, NOCl, O2, O3, OF2, P2, PF3, PH3, S2, SF2, SiH4, SiO2, and SO2.
-The performance of the recommended model parameters was assessed on the standard G2 set of molecules92, containing 118 closed- and 30 open-shell systems.
+The performance of the recommended model parameters was assessed on the standard G2 set of molecules97, containing 118 closed- and 30 open-shell systems.
 For training and assessment computations cc-pVXZ38, cc-pwCVXZ40(X = D, T, Q) and def2-SVP, def2-XZVP (X = T, Q)44 were used.
@@
 The MADF model parameters for use with relativistic OBS were determined on training set TS2 composed of 24 closed-shell systems with heavy elements: AgH, AsH3, AuH, BiH3, CuH, GaH3, GeH4, HAt, HBr, Hi, InH3, Kr, PbH4, PoH2, PtC, RuC, SbH3, ScH, SeH2, SnH4, TeH2, TlH3, Xe and Rn.
-The performance of the recommended model parameters was assessed on 2 sets: (a) the Ln54 benchmark set composed of 54 rare-earth compounds93, and (b) the set of 60 diatomics containing d-block (transition metal) atoms from Ref.
+The performance of the recommended model parameters was assessed on 2 sets: (a) the Ln54 benchmark set composed of 54 rare-earth compounds98, and (b) the set of 60 diatomics containing d-block (transition metal) atoms from Ref.
 99 (we will refer to this set as Tm60).
@@
 For cc-pVXZ, cc-pwCVXZ, and def2- family of OBSs we also compared against the corresponding manually-optimized DFBS cc-pVXZ-RIFIT, cc-pwCVXZ-RIFIT (X = D, T, Q) and def2-SVP-RIFIT and def2-XZVP-RIFIT (X = T, Q) for the G2 set for comparison.
-All geometries, except for the G2 set, were optimized with the PBE095 and def2-TZVP basis set using the Psi4’s96 geometry optimization module.
-For elements with atomic number greater than 36, the effective core potentials (ECPs)97 available for def2-TZVP were used for geometry optimization.
-All matrix pseudo-inverses and pseudo-inverse square roots were computed with Löwdin’s orthogonalization procedure98.
+acCD calculations were performed using OpenMolcas version 24.10100.
+Since this version of OpenMolcas does not have an implementation of open-shell MP2, for comparison of MADF and acCD only the closed-shell subset of G2 set was used.
+We also show acCD vs MADF comparison only for the cc-pVXZ family of basis sets.
+All geometries, except for the G2 set, were optimized with the PBE0101 and def2-TZVP basis set using the Psi4’s102 geometry optimization module.
+For elements with atomic number greater than 36, the effective core potentials (ECPs)103 available for def2-TZVP were used for geometry optimization.
+All matrix pseudo-inverses and pseudo-inverse square roots were computed with Löwdin’s orthogonalization procedure104.
 4 Results
 4.1 Training MADF model parameters
-The MADF model parameters were tuned to ensure that that the DF errors for the respective training sets did not exceed target accuracies of ±$\pm 20\mu E_{h}$ and ±$\pm 10\mu E_{h}$ per electron for HF and MP2 energies, respectively.
+The MADF model parameters were tuned to ensure that that the DF errors for the respective training sets did not exceed target accuracies of $\pm 20\mu E_{h}$ and $\pm 10\mu E_{h}$ per electron for HF and MP2 energies, respectively.
 This target range for errors was determined by analyzing the DF errors of manually-optimized DFBSs; the majority have larger errors than these targets (see Section 4.2.1).
-4.1.1 Training $\tau_{\text{H}}$
-The subset of all hydrogen-containing molecules from the TS1 set (Section 3) was used to determine the recommended value of $\tau_{\text{H}}$.
-While varying $\tau_{\text{H}}$ the remaining MADF parameters were fixed ($\tau_{1}=\tau_{2}=10^{-8},\zeta=1.2$) sufficiently closely to their asymptotic limits ($\tau\to 0$, $\zeta\to 1$) without causing ill-conditioning and excessive costs.
-Figures 2(a) and 2(b) show the DF errors of HF and MP2 energies, respectively, with $\tau_{\text{H}}$ varying between $10^{-8}$ and $10^{-4}$.
-The errors are essentially constant with $\tau_{\text{H}}\leq 10^{-6}$, thus $\tau_{\text{H}}=10^{-6}$ is recommended.
-Since hydrogen has almost negligible relativistic effects, this value will be used henceforth for both relativistic and nonrelativistic calculations.
-Figure 2: Variation of DF errors of nonrelativistic (a) HF and (b) MP2 energies of the H-containing subset of the TS1 training set vs the $\tau_{\text{H}}$ model parameter of the MADF generator.
 4.1.1 Training $\tau_{1,2}$ and $\zeta$ for nonrelativistic OBS
 The variation of DF errors with $\tau_{2}$ was studied next, by keeping the rest of the parameters fixed ($\tau_{1}=10^{-8},\zeta=1.2$) sufficiently closely to their asymptotic limits ($\tau\to 0$, $\zeta\to 1$) without causing ill-conditioning and excessive costs.
-As Figures 3(a) and 3(b) indicate, the DF errors of HF and MP2 are largely converged at $\tau_{2}=10^{-5}$.
-Next, $\tau_{2}$ was fixed at $10^{-5}$ and $\tau_{1}$ was varied (Figures 4(a) and 4(b)).
+As Figures 2(a) and 2(b) indicate, the DF errors of HF and MP2 are largely converged at $\tau_{2}=10^{-5}$.
+Next, $\tau_{2}$ was fixed at $10^{-5}$ and $\tau_{1}$ was varied (Figures 3(a) and 3(b)).
 The DF errors are sufficiently converged with $\tau_{1}=10^{-6}$.
-Finally, with $\tau_{1}=10^{-6}$ and $\tau_{2}=10^{-5}$ variation of DF errors with $\zeta$ was studied (Figures 5(a) and 5(b)).
+Finally, with $\tau_{1}=10^{-6}$ and $\tau_{2}=10^{-5}$ variation of DF errors with $\zeta$ was studied (Figures 4(a) and 4(b)).
 For $\zeta>1.4$ the max DF errors increase due to the insufficient spanning of the OBS AO product space by the DFBS AOs, whereas for $\zeta<1.2$ (not shown) the onset of ill-conditioning in DFBS makes the MADF model not numerically stable.
 $\zeta=1.4$ ensures sufficient convergence of the DF errors without undue numerical problems.
-Note that $\zeta=1.4$ is significantly smaller than the exponent ratio 1.8 used by the AutoAux generator ( 1.8)61 and ratio 2 used by PySCF’s generator65 for even-tempered DFBS construction.
+Note that $\zeta=1.4$ is significantly smaller than the exponent ratio 1.8 used by the AutoAux generator ( 1.8)61 and ratio 2 used by PySCF’s generator67 for even-tempered DFBS construction.
 Our findings confirm that the relatively small exponent ratios are indeed necessary for accurate density fitting.
@@
 Parameters $\tau_{1,2}$ and $\zeta$ were trained again using the relativistic HF and MP2 energies of the TS2 set.
-$\tau_{2}$ was varied first, with $\tau_{1}=10^{-8}$ and $\zeta=1.2$ (Figures 6(a) and 6(b)); just as in nonrelativistic computations $\tau_{2}=10^{-5}$ ensures sufficient convergence of the DF errors.
-$\tau_{1}$ was then varied with fixed $\tau_{2}=10^{-5}$ and $\zeta=1.2$ (Figures 7(a) and 7(b)).
+$\tau_{2}$ was varied first, with $\tau_{1}=10^{-8}$ and $\zeta=1.3$.
+Here we start with $\zeta=1.3$ instead of $\zeta=1.2$ like in the non-relativistic case because $\zeta=1.2$ is too small and causes numerical issues with the relativistic OBSs.
+(Figures 5(a) and 5(b)); just as in nonrelativistic computations $\tau_{2}=10^{-5}$ ensures sufficient convergence of the DF errors.
+$\tau_{1}$ was then varied with fixed $\tau_{2}=10^{-5}$ and $\zeta=1.3$ (Figures 6(a) and 6(b)).
 In contrast to the nonrelativistic case, a much smaller $\tau_{1}$ is needed for sufficient convergence of the DF errors.
-The variation of relativistic DF errors with $\zeta$ (Figures 8(a) and 8(b)) is similar to the nonrelativistic case, with $\zeta=1.4$ deemed sufficient.
-Yet again, this is a much smaller ratio than the ratios used by even-tempered DFBS generators.61, 65
+The variation of relativistic DF errors with $\zeta$ (Figures 7(a) and 7(b)) is similar to the nonrelativistic case, with $\zeta=1.4$ deemed sufficient.
+Yet again, this is a much smaller ratio than the ratios used by even-tempered DFBS generators.61, 67
 The optimal values of the parameters MADF are briefly tabulated in Table 1.
@@
 (c) Comparison of DFBS sizes, relative to that of the corresponding OBS.
+We have also compared the performance of primitive DFBSs produced by MADF with contracted DFBSs produced by acCD.
+From Figure 9(a) it is clear that MADF DFBSs produce smaller DF errors in HF energies compared to acCD DFBSs.
+As can be seen in Figure 9(b) the MP2 errors for cc-pVDZ are smaller with MADF compared to acCD and for cc-pVTZ and cc-pVQZ the magnitude MP2 errors are $5\mu E_{h}$ per electron, which is sufficient.
+MADF also provides significantly smaller DFBSs than that of acCD with virtually similar errors for cc-pVTZ and cc-pVQZ as seen in Figure 9(c).
+Figure 9: Comparison of DF errors of (a) HF and (b) MP2 energies of the closed-shell subset of the G2 test set obtained with acCD and MADF DFBSs.
+(c) Comparison of DFBS sizes, relative to that of the corresponding OBS.
+For acCD, the number of contracted DF AOs was used.
 4.2.2 Ln54 and Tm60 sets
@@
 This work introduced the MADF algorithm for generating density-fitting basis set composed of primitive solid-harmonic Gaussian AOs and suited for nonrelativistic and relativistic (all-electron) computations with mean-field and correlated electronic structure models.
-Not using even-tempered AO sets for spanning the AO product space, as in other comparable generators,61, 65 produces compact and accurate candidate pools of DFBS AOs while sufficiently reducing the numerical redundancy.
+Not using even-tempered AO sets for spanning the AO product space, as in other comparable generators,61, 67 produces compact and accurate candidate pools of DFBS AOs while sufficiently reducing the numerical redundancy.
 The subsequent pruning of the candidate DFBS pool based on the 2-body energy as the importance metric allows to naturally reduce the relevant range of angular momentum channels and exponent spans with minimal use of heuristics.
 Generation of primitive (rather than contracted) DF AOs allows to keep the integral evaluation costs optimally low.
-Using only 4 model parameters MADF generates basis sets that match or exceed the accuracy-to-size ratio of the state-of-the-art DFBS generators; the comparison is particularly favorable with OBSs suited for relativistic all-electron computations.
-Computations utilizing MADF DFBS, of course, can also benefit from system-specific DFBS compressions.67, 99
+Using only 3 model parameters MADF generates basis sets that match or exceed the accuracy-to-size ratio of the state-of-the-art DFBS generators; the comparison is particularly favorable with OBSs suited for relativistic all-electron computations.
+Computations utilizing MADF DFBS, of course, can also benefit from system-specific DFBS compressions.70, 105
 Although picture change and other relativistic effects on the 2-particle interaction were not considered here, it will be interesting to examine the impact of such effects on the optimal MADF model parameters in the future; it is known that there are nontrivial differences in the requirements on the DFBS that stem from relativistic contributions to the effective 2-particle interactions.13 Even in the nonrelativistic framework much more testing of the approach will be clearly needed.
@@
 Early experiences with MADF DFBSs appear promising.
-For example, we recently utilized the MADF DFBS generator for explicitly-correlated (F12) coupled-cluster computations with aug-cc-pV7Z OBS (for which no manually-optimized DFBS is available) where we found that the DF errors with the MADF DFBS are significantly smaller than the use of the manually-optimized aug-cc-pV6Z-RIFIT basis.
-These and other experiences with MADF DFBSs will be reported elsewhere.
+For example, we recently utilized the MADF DFBS generator for explicitly-correlated (F12) coupled-cluster computations with aug-cc-pV7Z OBS (for which no manually-optimized DFBS is available) where we found that the DF errors with the MADF DFBS are significantly smaller than the use of the manually-optimized aug-cc-pV6Z-RIFIT basis.106 These and other experiences with MADF DFBSs will be reported elsewhere.
 {suppinfo}
@@
 {acknowledgement}
-This material is based upon work supported by the U.S.
+Work by KAS was supported by the U.S.
 Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, and Office of Basic Energy Sciences, Scientific Discovery through the Advanced Computing (SciDAC) program under Award DE-SC0022263.
+Work by EFV was supported by the US Department of Energy, Office of Science, via award DE-SC0022327.
 This project used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S.
@@
 Appendix A Superposition of atomic densities (SOAD)
-Unlike the SAD method,85 with computes the Fock matrix from the atomic (optionally, spin-free) ensemble density matrix computed in OBS self-consistently, SOAD Fock matrix is computed from fixed spin-free ensemble occupancies ${\bf n}_{\mathrm{SOAD}}$ of the minimal basis set (MBS) AOs, without self-consistency.
+Unlike the SAD method,89 with computes the Fock matrix from the atomic (optionally, spin-free) ensemble density matrix computed in OBS self-consistently, SOAD Fock matrix is computed from fixed spin-free ensemble occupancies ${\bf n}_{\mathrm{SOAD}}$ of the minimal basis set (MBS) AOs, without self-consistency.
 ${\bf n}_{\mathrm{SOAD}}$ for a single atom in an orthonormal MBS is obtained by smearing each subshell’s electrons evenly among all of its orbitals; e.g., for a carbon atom occupancies of 1s, 2s, 2px, 2py, and 2pz MBS AOs are 2, 2, 2/3, 2/3, and 2/3, respectively.
-This recipe has been referred to as average-of-configuration (AOC)100 and canonical ensemble101.
-The number of electrons in each subshell is defined by the electron configuration of that atom’s neutral ground state taken from NIST Atomic Spectra Database102.
+This recipe has been referred to as average-of-configuration (AOC)107 and canonical ensemble108.
+The number of electrons in each subshell is defined by the electron configuration of that atom’s neutral ground state taken from NIST Atomic Spectra Database109.
 For a molecule ${\bf n}_{\mathrm{SOAD}}$ is obtained by concatenation of the atomic occupancies, without taking into account the effects of nonunit molecular MBS metric.
-In this work the MINI basis103, 104 was used as MBS for $Z\leq 54$, and the ANO-RCC-MB basis105 for $54<Z\leq 96$.
+In this work the MINI basis110, 111 was used as MBS for $Z\leq 54$, and the ANO-RCC-MB basis112 for $54<Z\leq 96$.
 ${\bf n}_{\mathrm{SOAD}}$ is used to construct SOAD Fock matrix in OBS AO basis, $\mathbf{F_{\mathrm{SOAD}}}$, as follows:
 $(\mathbf{F}_{\mathrm{SOAD}})^{\mu}_{\nu}=h^{\mu}_{\nu}+\sum_{m}({\bf n}_{\mathrm{SOAD}})_{m}[2(\mu\nu|mm)-(\mu m|\nu m)],$ (26)
-where $h^{\mu}_{\nu}$ are the matrix elements of the nonrelativistic 1-body (core) Hamiltonian, $m$ runs over MBS AOs and $\mu,\nu$ over OBS AOs.
+where $h^{\mu}_{\nu}$ are the matrix elements of the 1-body (core) Hamiltonian (sf-1eX2C90, 91, 92 in this work) , $m$ runs over MBS AOs and $\mu,\nu$ over OBS AOs.
+It should be noted that Eq.
+26 is similar to the projection-free guess proposed by Norman and Jensen113 for Dirac-Hartree-Fock/Kohn-Sham.